### v1

# HuggingGraph: Understanding the Supply Chain of LLM Ecosystem

Abstract.
Large language models (LLMs) leverage deep learning architectures to process and predict sequences of words based on context, enabling them to perform a wide range of natural language processing tasks, such as translation, summarization, question answering, and content generation. However, the increasing size and complexity of developing, training, and deploying cutting-edge LLMs demand extensive computational resources and large-scale datasets. This creates a significant barrier for researchers and practitioners. Because of that, platforms that host models and datasets have gained widespread popularity. For example, on one of the most popular platforms, i.e., Hugging Face, there are more than 1.8 million models and more than 450K datasets by the end of Jun. 2025, and the trend does not show any slowdown.
As existing LLMs are often built from base models or other pre-trained models and using external datasets, they can inevitably inherit vulnerabilities, biases, or malicious components that exist in previous models or datasets. Therefore, it is critical to understand these components’ origin and development process to detect potential risks better, improve model fairness, and ensure compliance with regulatory frameworks. Motivated by that, this project aims to study such relationships between models and datasets, which are the central parts of the LLM supply chain. First, we design a methodology to collect LLMs’ supply chain information systematically. With the collected information, we design a new graph to model the relationships between models and datasets, which is a large directed heterogeneous graph, with 397,376 nodes and 453,469 edges.
Then, on top of this graph, we perform different types of analysis and make multiple interesting findings, such as (i) the LLM supply chain graph is large, sparse, and exhibits a power-law degree distribution; (ii) it features a densely connected core and a fragmented periphery; (iii) datasets are pivotal, playing critical roles in training; (iv) a strong interdependence exists between models and datasets; (v) the graph is dynamic, with daily updates capturing the ecosystem’s continuous evolution.
1. Introduction
Large language models (LLMs) are AI models designed to understand and generate human language by learning patterns and relationships within extensive datasets (Mirchandani et al., 2023; Shen et al., 2023), such as GPT (Generative Pre-trained Transformer) (Yenduri et al., 2024), BERT (Bidirectional Encoder Representations from Transformers) (Kenton and Toutanova, 2019), and T5 (Text-To-Text Transfer Transformer) (Bahani et al., 2023). These models leverage deep learning architectures to process and predict sequences of words based on context, enabling them to perform a wide range of natural language processing tasks (Bonner et al., 2023) such as translation (Lu et al., 2024b), summarization (Laban et al., 2023), question-answering (Allemang and Sequeda, 2024), and content generation (Agossah et al., 2023). LLMs usually have billions (or even trillions) of parameters (Merrick et al., 2024), enabling them to generate high-quality text.
However, the increasing size and complexity of developing, training, and deploying cutting-edge LLMs demand extensive computational resources (Xu et al., 2024b) and large-scale datasets (Villalobos et al., 2024). This creates a significant barrier for researchers and practitioners, limiting their access to state-of-the-art models (Mizrahi et al., 2024). As the demand for democratizing access to such LLM models continues to rise, platforms that host models and datasets have gained widespread popularity. For example, Figure 1 shows the number of models and datasets (in a million scales) on Hugging Face, a popular AI model hosting platform (Face, [n. d.]), starting from July 31st, 2024, to Jun. 30th, 2025. By the end of June 2025, it has reached over 1.8M models and 450K datasets. In addition, the trend does not show any slowdown. Such LLM hosting platforms provide user-friendly interfaces, APIs, and cloud-based infrastructures that enable researchers and developers to easily share, fine-tune, and deploy models without requiring extensive computational resources. Moreover, they foster open collaboration, allowing the broader community to contribute to model improvements, benchmark performances, and enhance transparency in AI research.
On such platforms, different types of models can be classified into two categories based on their tasks, i.e., base models and task-specific models. (i) base models are large, pre-trained models that can be fine-tuned for specific downstream tasks (Tan et al., 2024). They are usually trained on vast datasets and are general-purpose, such as GPT (Yenduri et al., 2024), BERT (Kenton and Toutanova, 2019), and T5 (Bahani et al., 2023). (ii) Task-specific models are modified versions of base models for a specific task. Taking Hugging Face as an example, there are four types of such models. First, the fine-tuned models adapt base models for specific tasks by training on additional task-specific datasets (Zou et al., 2023). Second, the adapters models add lightweight and modular layers to the pre-trained models for specific tasks (Hu et al., 2023). Third, the quantization models trade off the precision in numerical computations for accelerating inference and reducing memory consumption (e.g., using less precise model parameters) (Velingker et al., [n. d.]). Fourth, the merge models integrate multiple specialized or fine‑tuned parameter sets into a single unified model by combining weights or configurations (e.g., via weighted averaging or parameter stitching), enabling support for multiple tasks or domains without separate deployments (Akiba et al., 2025). Besides models, such platforms also host many datasets used for training and fine-tuning the previously discussed models (Röttger et al., 2024).
Figure 1. The number of models and datasets (in millions) on Hugging Face from July 31, 2024, to June 30, 2025.
1.1. Motivation
As existing LLMs are often built from base models or other pre-trained models and using external datasets, they can inevitably inherit vulnerabilities, biases, or malicious components that exist in previous models or datasets. Therefore, understanding these components’ origin and development process can help better detect potential risks, improve model fairness, and ensure compliance with regulatory frameworks.
Motivated by that, this project aims to study such relationships between models and datasets. They are the central parts of the LLM supply chain (Wang et al., 2024), which refers to the entire lifecycle of developing, training, and deploying LLMs, similar to a traditional supply chain in manufacturing or software development (Čolaković et al., 2021; Anonymous, 2022; Chou et al., 2006; Sonatype, 2015). The LLM supply chain can help to identify critical insights for both model evolution and dataset origin, as discussed below.
Model evolution. The study of the supply chain of LLMs gives a clear overview of how LLMs evolve from base models to fine-tuned variants, adapter integration, quantization models, and merged models that combine multiple specialized parameter sets into a single release. With that, one can easily keep track of them. For example, a use case is when a security or bias risk is found in one LLM, and we can quickly locate the potential models that might have the same problems.
Dataset origin. This supply chain can help to understand the datasets’ origins used for training different models (Shen et al., 2023). Dataset origin refers to the source from which the data is collected. For example, for a fine-tuned model, we not only care about which dataset is used for fine-tuning but also what other datasets are involved in training the previous model. Understanding such dataset origin helps to ensure that the dataset used is reliable, ethical, legally compliant, and relevant to the task.
1.2. Contribution
Our contributions are mainly threefold. First, we design a methodology to systematically collect the supply chain information of LLMs. In this paper, we mainly study the most popular ML platform, i.e., Hugging Face, but the same strategy also applies to other LLM platforms. In particular, we use the APIs from the ML platform to collect the metadata about the hosted model and dataset. To that end, we collected a large dataset, covering all the models (1.8M) and datasets (450K) as of June 30, 2025.
Second, with the collected metadata, we designed a new graph, named LLM supply chain graph, to model the relationships between models and datasets. It is a directed heterogeneous graph where a node denotes different types of datasets and models (including base, fine-tune, adapter, quantization, and merge models). An edge denotes the dependency relationship between them, including dataset-dataset, and model-dataset relationships. Together, this graph is able to accurately capture the LLM supply chain information. To this end, we constructed a large graph with 397,376 nodes and 453,469 edges. An anonymized version of the complete graph is publicly available at GitHub [^1] Anonymized LLM supply chain graph: https://github.com/huggingface00/HuggingGraph.
Third, with this graph, we perform different types of analysis, including model-dataset relationship analysis. We answer five research questions, including (i) the properties of the LLM supply chain graph, (ii) the structural analysis, (iii) the supply chain relationships between datasets, (iv) the supply chain relationships between the models and datasets, and (v) $\Delta$-based temporal update and tracking.
2. Preliminary
The LLM supply chain encompasses the interconnected processes required for developing, deploying, and maintaining models (Wang et al., 2024). This includes sourcing and preparing data to ensure high-quality and diverse datasets (Wang et al., 2024). It also involves creating and training models (Li et al., 2023). Finally, it covers making trained models available through APIs (Singla et al., 2023). In addition, LLMs can undergo adaptation, quantization, and fine-tuning, a process where they are tuned with domain-specific datasets to maximize performance on specific tasks (VM et al., 2024), thus improving their accuracy and applicability.
Note that this study mainly focuses on the relationships between various types of models and datasets, which are the central parts of the whole LLM supply chain ecosystem. We hope this study can not only provide insights on the relationships between LLM models and datasets but also raise awareness and future research interests in the LLM supply chain ecosystem.
3. Methodology
This section discusses the main methodologies, including data collection, LLM supply chain graph construction, and analysis.
Table 1. The APIs used to extract data from Hugging Face.
API Name Description
Model hub API Access the model hub to list, search, and download models and metadata.
Dataset API Access the datasets for discovery, metadata retrieval, and downloading.
Metrics API Access metrics for model evaluation, e.g., metric discovery, metadata retrieval, and calculation.
Search API Search the name, tag, or other metadata for models and datasets.
3.1. LLM Supply Chain Information Collection
To analyze the LLM supply chain ecosystem, we need a large dataset with such information. Fortunately, platforms like Hugging Face provide some APIs that allow us to access the model and dataset and collect their metadata information, which can be used to construct the supply chain ecosystem.
Table 1 summarizes four types of APIs we used. In particular, (i) the model hub APIs allow one to access the hub of existing models, including searching and downloading the model and its metadata. (ii) The dataset APIs allow one to access the datasets for discovery, metadata retrieval, and downloading. (iii) The metrics APIs allow one to access the metrics for model evaluation, including metric discovery, metadata retrieval, and calculation. (iv) In addition, one can use search APIs to search the name, tag, or other metadata for models and datasets.
With the help of these APIs, we are able to collect the metadata of the hosted models and datasets. In particular, we collected a large dataset, covering the models (330K) and datasets (66K) as of June 30, 2025.
Handling missing information. Accurate construction of the LLM supply chain graph critically depends on the quality of metadata from the LLM platforms (e.g., Hugging Face), which might suffer from missing or incomplete data. To address this limitation, we design the following two techniques, i.e., cross-reference links and textual pattern extraction.
(i) Cross-reference links. The model or dataset description could miss the supply chain data fields for API queries, which could be embedded within statically or dynamically rendered HTML pages. In this example URL https://huggingface.co/models?other=base_model:finetune:meta-llama/Meta-Llama-3-8B, one can tell the model “Meta-Llama-3-8B” is fine-tuned from the model in the previous webpage. To capture such information, we cross-reference the links of the filtered model listing webpages, extract model identifiers, enable the reconstruction of supply chain graph edges, and recursively trace model lineage from the leaf node. This scraping step complements API-based extraction and is only employed when reverse dependency data is otherwise inaccessible.
Figure 2. An example subgraph centering on the base model “Meta-llama” from the complete LLM supply chain graph.
(ii) Textual pattern extraction. When structured metadata is absent, the model and dataset cards might mention dependencies in unstructured text descriptions. To capture that, we employ a named entity recognition (NER) method (Vasiliev, 2020; Li et al., 2020) to extract the dependency relationships from the text. For example, the textual phrase like “fine-tuned from Llama-2” contains the fine-tuned keyword, which implies from which model this model is actually fine-tuned. Similarly, we also look into other words, such as “train”, and “adapt”.
3.2. LLM Supply Chain Graph
With the collected metadata, we construct a directed heterogeneous graph to model the LLM supply chain. In this graph, a node denotes different types of datasets and models, including base, fine-tuned, adapter, and quantization models.
An edge denotes the dependency relationship between them, including model-model, dataset-dataset, and model-dataset relationships. Figure 2 shows a simplified supply chain subgraph centering on the base model Meta-llama.
(i) Dataset-dataset relationship. The datasets within the LLM supply chain might overlap, build upon, or extend from each other–e.g., a dataset may be a subset or modified version of another. To capture such information, we connect them with two types of edges. (1) Subset relationship. These arise when a dataset is explicitly documented as a subsample or partition of another. For instance, a dataset named “C4_200M” is described as a subset of “C4”. (2) Modified version. These represent updates or enhanced variants of existing datasets. For example, “TruthfulQA_v2” incorporates factual corrections and improvements over an earlier version, “TruthfulQA_v1”.
(ii) Model-dataset relationship. To capture the model and dataset dependency, we use the metadata information from both models and datasets. From the metadata of a model, it might include the datasets used to train it. However, not all the models disclose such information. To find more information, we observe the metadata of a dataset, including a data field on “trained_fine_tune_models” on this dataset. Thanks to that, we are able to capture the accurate model and dataset relationship. In Figure 2, the datasets The Pile and Awesome-Chatgpt-prompt have directed edges to model Meta-Llama, meaning that both datasets are used to train the model.
3.3. Supply Chain Graph Analysis
This supply chain graph can help to understand the transformational processes of the models and datasets. In particular, we can understand how base models evolve into their variants, including fine-tuned, adaptive, and quantized models, and vice versa. Similar observations can be made for datasets. This would provide a clear view of how the base model (or dataset) is transformed for performing a particular task. In particular, we mainly perform two types of analysis, i.e., forward and backward analysis.
Dataset analysis example. For the dataset, our supply chain analysis shows how different datasets connect and form a new dataset and work together. This combination creates flexible resources that improve how models perform in various areas. In Figure 2, the dataset The Pile is composed of multiple subsets, including Wikimedia, Arxiv, Openwebtext2, and Pubmed Central. Together, these datasets form a unified corpus that serves as training data for models like Meta-LlaMA.
Model and dataset analysis example. In Figure 2, to analyze the backward supply chain of the model RBot70Bv4, we trace its lineage back through its development stages. This model is fine-tuned from Unsloth, which in turn originates from its base model, Meta-Llama, and the datasets used to train the base model are The Pile and Awesome-Chatgpt-prompts. Through this analysis, we establish the backward path, starting from the target model—RBot70Bv4—and tracing back to its base model, Meta-Llama, revealing the dependencies and transformations involved in its development. Similarly, we can analyze the datasets.
3.4. Temporal Update and Tracking
To ensure HuggingGraph remains an accurate and evolving representation of the LLM ecosystem, we have implemented a (i) $\Delta$-based batch update mechanism and (ii) a versioned archival system.
$\Delta$-based graph update. To ensure HuggingGraph reflects the evolving nature of the LLM ecosystem, we implement a $\Delta$-based batch update mechanism. Instead of reconstructing the graph from scratch, we compute incremental updates by comparing successive versions. Let $G_{t}$ denote the graph at time $t$, and define the $\Delta$ as $\Delta_{t}=G_{t}-G_{t-1}$, where $\Delta_{t}$ captures all added or removed nodes and edges. This approach offers both computational scalability and storage efficiency. At scheduled intervals (e.g., daily), we collect metadata from Hugging Face and compute the differences. For instance, between June 25–July 15, 2025, the platform added 104,380 new models ($\sim$3,866/day), 20,034 new datasets ($\sim$742/day), and approximately 854 dependency edges/day.
Versioning and longitudinal tracking. To enable temporal and historical analyses, each update generates a versioned snapshot of the graph, archived with a timestamp and associated change log. This log captures additions, deletions, and dependency modifications, such as new edges from TruthfulQA v2 to TruthfulQA v1. This ensures that HuggingGraph remains a living resource that adapts to ongoing developments in LLM research and deployment.
4. Experiment and Finding
To deeply understand the relationships between models and datasets, we aim to answer seven critical research questions (RQs). We believe they can offer valuable insights into the supply chain of the LLM ecosystem and potentially pave the way for future LLM development and deployment.
  • RQ #1: What are the properties of the LLM supply chain graph?
  • RQ #2: What structural patterns emerge from the LLM ecosystem?
  • RQ #3: What are the supply chain relationships between datasets?
  • RQ #4: What are the relationships between models and datasets?
  • RQ #5: What insights can be gained as the graph temporally updating?
4.1. RQ #1: Supply Chain Graph Properties
Figure 3. (a) Indegree distribution and (b) Outdegree distribution. The X-axis represents the number of indegrees and outdegrees, respectively, and the Y-axis represents the number of vertices in a logarithmic scale.
This research question aims to understand the critical properties of the LLM supply chain graph, i.e., graph basics and degree distribution.
Graph basics. The collected supply chain graph is a large directed heterogeneous graph with 397,376 nodes and 453,469 edges as of June 30th, 2025. In particular, there are six different types of nodes, including 25,717 base models, 115,103 fine-tuned, 79,240 adapters, 98,125 quantization models, 13,027 merges and 66,164 datasets. The average degree is about 1.14, denoting that it is a very sparse graph.
The degree distribution of a graph describes how node degrees (the number of edges connected to a node) are distributed across the graph. Figure 3(a) and (b) illustrate the indegree and outdegree distribution of the graph, respectively. We show not only the total distribution but also the distribution of five types of nodes, including base models, fine-tuned models, adapter models, quantization models, and datasets. We made two interesting observations.
(i) This degree distribution in our supply chain graph shows a heavy-tailed behavior. In particular, the indegree distribution shows a large spread across different categories. The outdegree distribution follows a similar pattern but may differ in specific cases (e.g., adapters seem to have a more restricted degree distribution). The power-law behavior suggests that most nodes have low degrees, while a few central nodes (hubs) dominate the graph. In particular, the model images from macrocosm-os has the highest indegree value of 550, and Mistral-7B-v0.1 from mistralai has the highest outdegree value of 1,093. Specifically, the base models act as high-degree hub nodes as they are heavily used by other task-specific models. To this end, we can conclude that this graph is a power-law graph, also known as a scale-free graph. That is, a graph whose degree distribution follows a power law distribution. In other words, a small number of nodes have many connections, while most nodes have a few connections.
(ii) We observe that all the base models have at least one outdegree as they are used by the task-specific models. For the dataset nodes, there are 20,093 datasets with zero indegrees, meaning they are the minimum datasets that can no longer be partitioned.
Finding #1: The LLM supply chain graph is a large, sparse graph that exhibits a power-law degree distribution, where a small number of high-degree base models and datasets serve as central hubs.
4.2. RQ #2: Supply Chain Structural Analysis
This research question aims to understand the topology and evolution of the LLM supply chain. To achieve that, we analyze the structural properties with connectivity and community analysis.
Connectivity analysis in a graph is crucial for understanding how well different nodes are linked and helps identify critical nodes and edges that maintain overall connectivity. This graph is a directed acyclic graph (DAG) because there are no dependencies in a cycle form, which also means there are no strongly connected components. Because of that, we only perform weakly connected components (WCC) analysis, and the total number of WCCs in our supply chain graph is 44,417.
Figure 4. Cumulative distribution function (CDF) of WCC size.
Figure 4 shows the cumulative distribution function (CDF) of the WCC distribution. We made two interesting observations. (i) The largest WCC covers 244,063 nodes, accounting for 61.4% of all the nodes. It reflects the dense interconnections that pervade the ecosystem. This vital element is essential for effective information sharing, resource allocation, and structural support and is the base of the ecosystem. In the largest WCC, major models are included, such as Gemma-2B, DistilBERT, and GPT-2.
(ii) In contrast, the remaining 153,313 WCCs collectively hold 31.4% of the nodes, with most having 1, 2, to 3 nodes. This indicates a fragmented outer edge characterized by specialized models, less frequent datasets, or active experimental projects. The prevalence of these small, isolated pieces suggests niche attempts that lack integration with the overall system. For example, tblard/tf-allocine (French sentiment), distilbert-base-uncased-finetuned-sst-2-english (sentiment analysis), and PolyAI/minds14 (banking queries) remain disconnected due to limited reuse or insufficient metadata.
Community analysis is the process of identifying groups or clusters of nodes within a graph that are more densely connected internally than with the rest of the graph. In the context of the LLM supply chain, it reveals semantically or functionally aligned subgraphs that reflect the patterns of reuse and task-specific models or datasets. We use the Louvain method (De Meo et al., 2011), which is a greedy optimization method for maximizing modularity. Modularity is a measure of how well a graph is divided into communities, where a high modularity score indicates that most edges fall within communities rather than between them. It compares the actual density of edges inside communities to the expected density if edges were placed at random.
Table 2 summarizes the top-10 communities detected using the Louvain method. Our analysis reveals an exceptionally modular structure, with each of the top communities achieving a high modularity score of 0.97, indicative of strong intra-community connectivity. Collectively, these communities span a wide range of functional domains, underscoring the presence of well-defined, task-aligned clusters within the ecosystem.
We made two interesting observations. (i) The largest community (ID 1) consists of 9,158 nodes and attains a modularity score of 0.97, indicating extremely cohesive internal structure. It revolves around base models such as OLMoE and CausalLM, and general-purpose datasets like prompt-perfect. This suggests a densely connected nucleus facilitating widespread reuse and fine-tuning—effectively functioning as the backbone of the LLM ecosystem.
(ii) Several other communities reflect clear task-based segmentation. For instance, Community 3 ( 7,058 nodes, modularity 0.97) focuses on benchmarking, with models like Wanxiang and datasets such as marco-o1. Similarly, Community 6 (5,030 nodes, modularity 0.97) centers on instruction-tuning, connecting models like MedLlama-3-8B with curated datasets like dpo-mix.
Table 2. Top-10 Louvain communities in the HuggingGraph.
ID Size E.g. models E.g. datasets Modularity
1 9158 OLMoE, CausalLM prompt-perfect 0.97
2 7058 qwen2.5_math Marco-o1 0.97
3 6987 Wanxiang smartllama3.1 0.97
4 5696 tinyllama Llama-1B 0.97
5 5355 Qwen2.5-32B Matter-0.2 0.97
6 5030 MedLlama-3-8B dpo-mix 0.97
7 4297 Electra, ArliAI MixEval 0.97
8 4133 aesqwen1.5b llava 0.97
9 4067 bert TORGO 0.97
10 3932 Mistral vicuna_format 0.97
Finding #2: The LLM supply chain graph reveals a structure with a densely connected core and a fragmented periphery. A single dominant component encompasses the majority of nodes, supporting efficient model reuse and dependency tracing, while many smaller, isolated components reflect specialized or under-integrated efforts. Community analysis shows strong modularity, with nodes clustering around shared tasks and functional roles.
4.3. RQ #3: Dataset Supply Chain Analysis
This research question aims to provide comprehensive insights into the different usages of the datasets within the LLM supply chain graph. There are two relationships between datasets. On one hand, one dataset can be created by combining a variety of other datasets from disparate sources. On the other hand, one dataset can serve as a building block for new datasets, which enables continuous improvement in machine learning assets. In the following, we discuss the impacts on both training and benchmarking datasets.
Training dataset impact. In the LLM supply chain graph, 66,164 datasets act as training datasets that are used for model training or generating new datasets. Table 3 presents the top 10 datasets ranked by the highest number of included and derived datasets.
(i) We observe that a training dataset can include many small datasets. Sitting at the top is macrocosm-os/images (Kido, 1961) from macrocosm. It includes 550 sub-datasets and is a general-purpose image modification corpus used in advance multimodal AI research and large-scale vision-language model training. Moreover, bespokelabs/Bespoke (Elliot et al., 2024) from bespokelabs includes 215 sub-datasets, is a high-quality synthetic dataset designed to enhance multimodal AI research.
(ii) The other observation is that a single dataset can be included in many larger datasets, showing the huge overlaps between the datasets and also the models trained with them. The right two columns of Table 3 show the top 10 datasets sorted by the maximum number of derived datasets they have been included. The leading contributor, HuggingFaceH4 from HuggingFace has been involved in 989 derived datasets. It is primarily used for natural language understanding and generation tasks, including chatbots, text generation, code completion, reasoning, and multilingual processing, enabling the training and enhancement of models specialized in advanced conversational and instruction-following capabilities (Dietrich and Hollstein, 2025). Similarly, the CodeFeedback dataset from m-a-p has facilitated the creation of 857 derived datasets, is a collection of high-quality code instruction queries designed to enhance the training of large language models (LLMs) for code generation, debugging, and explanation tasks, enabling improved performance in complex programming scenarios(Meyers et al., 2018).
Table 3. Top-10 datasets sorted by the # of included and # of derived datasets.
Dataset
# of included
dataset
Dataset
# of derived
datasets
macrocosm-os/images 550 HuggingFaceH4 989
bespokelabs/Bespoke 215 CodeFeedback 857
databricks/databricks 137 MADLAD 850
Open-Orca/OpenOrca 119 Capybara 823
nguha/legalbench 117 Glot500 804
OpenAssistant/oasst1 111 dolphin-coder 756
LDJnr/Capybara 104 SlimOrca 726
kvn420/Tenro_V4.1 101 orca_dpo_pairs 687
teknium/OpenHermes 98 gutenberg 675
Anthropic/hh-rlhf 96 samantha 658
Finding #4: The datasets play critical roles in model training with 66,164 datasets are contributing to model development through inclusion and derivation, as seen in macrocosm-os/images and HuggingFaceH4.
4.4. RQ #4: Supply Chain between LLM Models and Datasets
This research question aims to explore the interconnections between models and datasets within the supply chain graph. It provides valuable insights from dual perspectives, as discussed below, including one dataset versus multiple models and one model versus multiple datasets.
One dataset versus multiple models refers to the case when a single dataset is used to train multiple models. Table LABEL:tab:dataset_model_1 shows the top 10 datasets based on the number of models trained on them. In particular, Mistral-7B-v0.1 takes the leading position and is a widely adopted open-source dataset known for its strong performance in general-purpose language understanding and generation tasks. It has been used to train 1,093 models, including 300 fine-tuned variants, 300 adapters, 193 quantized models, and 300 merged models, highlighting its broad adoption across diverse model derivation strategies.
In addition, the dataset TinyLlama-1.1B-v1.0, a compact and efficient model variant designed for low-resource deployment, is used to train 728 models, featuring 300 fine-tuned variants and 300 adapters. Similarly, open_llama_3b, an open-access dataset of LLaMA, supports 285 adapter-based models, indicating a preference for lightweight, modular adaptation. The dataset Yarn-Mistral-7b-128k also shows significant reuse, contributing to 279 adapters and 14 quantized models. Furthermore, MAmmoTH2-7B stands out with 210 merged models, showcasing its role in ensemble-style model fusion rather than traditional fine-tuning or adapter strategies. Lastly, the dataset Starling-LM-7B-alpha, known for alignment-focused training, contributes to a diverse range of downstream models, including 165 adapters and 18 quantized versions, illustrating its utility in fine-tuning and compression workflows.
One model versus multiple datasets refers to the case when an LLM model is trained with multiple datasets. Table LABEL:tab:dataset_model_2 shows the top 10 models ranked by the number of datasets used for training. DeBERTa-ST-AllLayers-v3.1, a fine-tuned variant of the DeBERTa architecture, takes the top position, having been trained on 116 different datasets. Its adapter-based counterpart, DeBERTa-ST-AllLayers-v3.1bis, also leverages the same number of datasets via adapter-based training, emphasizing modular reuse across tasks.
In addition, models like static-similarity-mrl-mul-v1 and static-similarity-mrl-multilingual are both fine-tuned on 108 datasets, indicating their role in multilingual and multi-task similarity-based retrieval applications. The ModernBERT-base-embed and Llama-3.2-3B-Instruct families show strong dataset diversity as well, with 88 and 87 datasets respectively, across fine-tuning and quantized variants (e.g., GGUF format). Interestingly, most of these models are fine-tuned—specifically, 8 out of the top 10—highlighting a trend where base models are transformed into task-specific variants through diverse training datasets. This pattern suggests that fine-tuning remains a dominant strategy for adapting base models to downstream tasks across heterogeneous data sources.
Finding #5: There exists a strong interdependence between models and datasets, where widely adopted datasets like Mistral-7B-v0.1 and TinyLlama-1.1B-v1.0 serve as the base for training hundreds of models across different adaptation techniques, while models such as DeBERTa-ST-AllLayers-v3.1 leverage diverse datasets to enhance their versatility, highlighting the critical role of dataset-model interactions in advancing AI capabilities.
Table 4. Top-10 datasets sorted by # of models trained.
Dataset
 Total
Fine-
tune
Adap-
ter
Quanti-
zation
Merges
Mistral-7B-v0.1 1093 300 300 193 300
TinyLlama-1.1B-v1.0 728 300 300 100 28
open_llama_3b 304 15 285 4 0
Yarn-Mistral-7b-128k 301 8 279 14 0
WizardVicuna-open-llama 280 12 261 7 0
TinyLlama-1.1B-v0.6 266 10 243 13 0
Yarn-Mistral-7b-64k 248 0 242 6 0
Nous-Capybara-7B-V1 213 11 174 27 1
MAmmoTH2-7B 213 0 0 3 210
Starling-LM-7B-alpha 210 10 165 18 17
Table 5. Top-10 models sorted by the # of datasets used for training.
Model
Model Type
# of dataset used-
for training
DeBERTa-ST-AllLayers-v3.1 Fine tune 116
DeBERTa-ST-AllLayers-v3.1bis Adapters 116
static-similarity-mrl-mul-v1 Fine tune 108
static-similarity-mrl-multilingual Fine tune 108
ModernBERT-base-embed Fine tune 88
Llama-3.2-3B-Instruct Fine tune 87
Llama-3.2-3B-Instruct-GGUF Quantization 87
DavidLanz-3.2-3B-Instruct Fine tune 87
static-retrieval-mrl-en-v1 Fine tune 79
XLMRoBERTaM3-CustomPoolin Fine tune 72
4.5. RQ #5: Temporal Update and Tracking
This research question aims to focus on the value of tracking fine-grained, time-based changes in the LLM supply chain. Our $\Delta$-based update mechanism (discussed in Section 3.4), produces a daily snapshot of the HuggingGraph, capturing precisely how many nodes and edges are added each day, how that translates into net growth, and what the busiest single day looks like. To better understand the implications of this tracking, we adopt both qualitative and quantitative perspectives on the evolving LLM ecosystem.
Fine-grained temporal insights into the LLM ecosystem. We would like to understand how the LLM supply chain evolves on a day-to-day basis. Hence, we adopt this fine-grained temporal qualitative approach that highlights fluctuations in model and dataset activity. This strategy allows us to uncover ecosystem-level dynamics such as release bursts, contributor behavior, and category-specific trends over time.
Figure 5 visualizes the daily $\Delta$-based growth of five key node categories (base, fine-tuned, adapters, quantized variants models, and datasets) from June 25 to July 15, 2025.
We observe that, (i) the fine-tuned models dominate the daily activity, averaging over 1,700 new entries per day, followed by consistent contributions from adapters ($\sim$750/day) and datasets ($\sim$700/day). Clear spikes on specific days—e.g., July 7 and July 9—coincide with major releases such as the Mistral-Fusion-v3 fine-tuning wave and dataset refreshes like HFTime2025-News. (ii) Furthermore, adapter uploads peaked at 887 on June 28, while quantized variants reached 290 on July 9, driven by releases such as QWin-GGUF-7B. These patterns demonstrate HuggingGraph’s ability to capture evolving supply chain dynamics at a fine-grained, time-sensitive resolution, revealing ecosystem bursts and contributor behaviors in real time.
Quantifying daily incremental graph growth. To complement our qualitative insights with measurable evidence, we adopt a quantitative approach that captures the scale and efficiency of daily changes to the LLM supply chain. This allows us to evaluate the performance of our $\Delta$-based update mechanism, ensuring support for large-scale, time-sensitive graph maintenance in a computationally efficient and auditable manner.
Table 6 provides a detailed breakdown of daily additions, net changes, and peak updates across node and edge types from Jun. 25 to Jul. 15, 2025. (i) On average, Hugging Face added 126 new fine-tuned models per day (peaking at 2,090), followed by adapters and datasets with 41 and 39 daily insertions, respectively. Quantized and merged models contributed 16 and 1 additions per day, respectively. In total, 2,526 fine-tuned models and 785 datasets were added during this period. (ii) Each day also saw approximately 23 $\pm$ 6 new dependency edges—linking models to their parent checkpoints or training datasets—consistent. By comparing successive graph versions, our update mechanism applies only the incremental changes ($\Delta_{t}$) rather than rebuilding the full graph—keeping updates both scalable and auditable in practice.
Figure 5. The temporal update for the datasets and different models on Hugging Face from June 25 to July 15, 2025.
Finding #6: The $\Delta$-based update mechanism offers clear visibility into the LLM ecosystem’s rapid and continuous evolution, with daily growth primarily driven by fine-tuned, adapter, and quantized models. It enables efficient, audit-friendly tracking by applying only incremental changes, eliminating the need to rebuild the entire graph. These capabilities ensure that HuggingGraph remains both scalable and precise in capturing the dynamic nature of the LLM supply chain.
5. Related Work
This section aims to provide in-depth coverage of the current research landscape concerning the opportunities within the LLM supply chain, mainly emphasizing the different dimensions of the relationship between the models and datasets, focusing on the model and dataset hosting platforms (Sindhu et al., 2024; Hadi et al., 2024).
LLM supply chain perspectives in AI. LLMs mark the revolutionary era in AI, with the edge passing through an exponential growth phase. A recent paper presents LLMs key components, including model infrastructure, lifecycle, and downstream applications (Wang et al., 2024). Another work showed that reusing these models has become so widespread, which encourages the sharing and adaptation of foundation models on a larger scale (Jiang et al., 2023). An open-source AI ecosystem, such as Hugging Face, hosts a broad range of LLMs and datasets and, therefore, plays an important role in democratizing AI technologies (Riva et al., 2024). The foundation models are key ingredients in such an ecosystem (Xu et al., 2024b). They can capture vast amounts of information from various datasets, allowing for effective task-specific variants like fine-tuning (Han et al., 2021). This trend underscores the significant impact of technological advancements on the democratization and innovation in AI (Cupać et al., 2024).
Table 6. Daily $\Delta$ statistics for HuggingGraph updates (June 25–July 15, 2025), showing average additions, net changes, and peak values by node and edge type.
Category
New-
per day
Net Change-
per day
Peak-
in month
Datasets 39 $\pm$ 62 +785 849
Base model 2 $\pm$ 5 +57 42
Fine-tune 126 $\pm$ 168 +2526 2090
Adapter 41 $\pm$ 64 +828 887
Quantized 16 $\pm$ 19 +321 290
Merge 1 $\pm$ 2 +34 13
Dependency edges 23 $\pm$ 6 +189 289
Relationship analysis between LLM models and datasets. A recent study investigates the practical adaptation of foundation models to specific domains and tasks. Multitask fine-tuning has demonstrated the potential to enhance performance on target tasks with scarce labels (Xu et al., 2024a). In plant phenotyping, adapting vision foundation models by techniques like adapter tuning and decoder tuning has shown results comparable to those of leading task-specific models (Chen et al., 2023). The Quadapter technique for language models tackles quantization difficulties by incorporating learnable parameters that scale activations channel-wise, mitigating overfitting during quantization-aware training (Park et al., 2022). The PathoTune framework in pathology imaging utilizes multi-modal prompt tuning to reconcile domain discrepancies among foundation models, tasks, and specific instances, exhibiting enhanced performance compared to single-modality methods and facilitating the direct adaptation of natural visual foundation models to pathological tasks (Lu et al., 2024a). These findings underscore the efficacy of various adaptation approaches in improving foundation model performance across multiple areas.
From the above, we can see that models heavily depend on each other for fine-tuning, adaptation, or quantization. Building on these insights, two other works have indicated that serious security vulnerabilities may encumber the LLM supply chain to a few developers (Kathikar et al., 2023; Osborne et al., 2024). These factors may affect the diversity of innovation. The authors relate several challenges arising during the software engineering, security, and privacy of different relations and components involved in model creation and deployment (Huang et al., 2023; Wang et al., 2024). Moreover, the vulnerabilities can be transferable from one model to another model during the fine-tuning process (Hajipour et al., 2024). Therefore, serious security challenges are also posed because the vulnerabilities of the repositories threaten the whole ecosystem (Kathikar et al., 2023).
Comparison with HuggingGraph. Our work is different from prior works in three aspects. First, the motivation for this work stems from the transferable property (Hajipour et al., 2024) of one model to others and during the training process with the datasets. Because of this, our analysis is mainly about the hierarchical relationship between the models, which model is used to fine-tune, adapt, or quantize other models, and which datasets are used to train the models. As it is challenging to get consistent hierarchical relational information, we made a supply chain graph from the metadata using APIs we got from the models and datasets. Second, as our analysis is one of the premier tasks in constructing a supply chain graph, for analyzing the relationship, we focus on the different graph properties like connectivity and WCC analysis. Third, we use forward and backward supply chain analysis to examine the built graph and learn more about the hierarchies of the models and datasets during fine-tuning, adaptation, quantization, and training. This is different from previous methods, which mostly focused on training the models. This would facilitate determining and identifying the relationship between the model and the dataset. More importantly, we conclude with multiple interesting findings, which we believe could provide important insights to the LLM community.
6. Discussion
HuggingGraph presents a technique to analyze the supply chain of the LLM ecosystem. The proposed graph can be used for various applications, e.g., auditing provenance, identifying biases, and revealing trends like quantized model scarcity, aiding researchers and platform maintainers. We discuss the following two use cases.
Use case #1: Tracing lineage and dependencies in the LLM supply chain. In the LLM ecosystem, models are frequently built upon others through fine-tuning, adapter training, or quantization, forming complex chains of dependencies. However, when these relationships are not explicitly visible, it becomes difficult to verify where a model comes from, whether it inherits bias from upstream datasets, or if it complies with licensing constraints. HuggingGraph can be used to address this challenge by constructing the supply chain of models and datasets, uncovering both direct and derived dependencies, even when they are not formally documented. For example, it can trace how the model Meta-llama indirectly relies on a dataset like Wikimedia via Awesome-Chatgpt-prompt (Figure 2). This transparency supports developers, auditors, and policymakers in validating provenance, detecting risks, and enabling responsible, trustworthy AI deployment.
Use case #2: Identifying critical nodes and structural vulnerabilities. In the LLM ecosystem, certain models (e.g., gemma-2b) and datasets (e.g., The Pile) are reused so frequently that they become critical structural hubs, where failure or removal of them could disrupt numerous downstream dependencies. These hidden single points of failure are difficult to detect without a comprehensive view of resource interconnections. HuggingGraph can be used to address this by modeling the supply chain as a graph and analyzing node connectivity to surface highly reused models and datasets with significant inbound or outbound links. This visibility enables maintainers to safeguard vital assets and helps developers mitigate the risk of overreliance on fragile or under-maintained components, fostering more robust and secure model development pipelines.
7. Conclusion
This project studies the relationships between models and datasets in the LLM ecosystem, which are the central parts of the LLM supply chain. First, we design a methodology to systematically collect the supply chain information of LLMs. With the collected information, we design a new graph to model their relationships, which is a large directed heterogeneous graph, having 397,376 nodes and 453,469 edges. From there, we perform different types of analysis and make multiple interesting findings.
Acknowledgment
This work was supported in part by National Science Foundation grants 2331301, 2508118, 2516003, 2419843. The views, opinions, and/or findings expressed in this material are those of the authors and should not be interpreted as representing the official views of the National Science Foundation, or the U.S. Government.

### v2

@@
 Because of that, platforms that host models and datasets have gained widespread popularity.
-For example, on one of the most popular platforms, i.e., Hugging Face, there are more than 1.8 million models and more than 450K datasets by the end of Jun.
-2025, and the trend does not show any slowdown.
-As existing LLMs are often built from base models or other pre-trained models and using external datasets, they can inevitably inherit vulnerabilities, biases, or malicious components that exist in previous models or datasets.
+For example, on one of the most popular platforms, i.e., Hugging Face, there are more than 1.8 million models and more than 450K datasets by the end of June 2025, and the trend does not show any slowdown.
+As existing LLMs are often built from base models or other pre-trained models and use external datasets, they can inevitably inherit vulnerabilities, biases, or malicious components that exist in previous models or datasets.
 Therefore, it is critical to understand these components’ origin and development process to detect potential risks better, improve model fairness, and ensure compliance with regulatory frameworks.
@@
 First, we design a methodology to collect LLMs’ supply chain information systematically.
-With the collected information, we design a new graph to model the relationships between models and datasets, which is a large directed heterogeneous graph, with 397,376 nodes and 453,469 edges.
-Then, on top of this graph, we perform different types of analysis and make multiple interesting findings, such as (i) the LLM supply chain graph is large, sparse, and exhibits a power-law degree distribution; (ii) it features a densely connected core and a fragmented periphery; (iii) datasets are pivotal, playing critical roles in training; (iv) a strong interdependence exists between models and datasets; (v) the graph is dynamic, with daily updates capturing the ecosystem’s continuous evolution.
+With the collected information, we design a new graph to model the relationships between models and datasets, which is a large directed heterogeneous graph, having 402,654 nodes and 462,524 edges.
+Then, on top of this graph, we perform different types of analysis and make multiple interesting findings.
 1. Introduction
 Large language models (LLMs) are AI models designed to understand and generate human language by learning patterns and relationships within extensive datasets (Mirchandani et al., 2023; Shen et al., 2023), such as GPT (Generative Pre-trained Transformer) (Yenduri et al., 2024), BERT (Bidirectional Encoder Representations from Transformers) (Kenton and Toutanova, 2019), and T5 (Text-To-Text Transfer Transformer) (Bahani et al., 2023).
-These models leverage deep learning architectures to process and predict sequences of words based on context, enabling them to perform a wide range of natural language processing tasks (Bonner et al., 2023) such as translation (Lu et al., 2024b), summarization (Laban et al., 2023), question-answering (Allemang and Sequeda, 2024), and content generation (Agossah et al., 2023).
+These models leverage deep learning architectures to process and predict sequences of words based on context, enabling them to perform a wide range of natural language processing tasks (Bonner et al., 2023), such as, translation (Lu et al., 2024), summarization (Laban et al., 2023), question-answering (Allemang and Sequeda, 2024), and content generation (Agossah et al., 2023).
 LLMs usually have billions (or even trillions) of parameters (Merrick et al., 2024), enabling them to generate high-quality text.
@@
 As the demand for democratizing access to such LLM models continues to rise, platforms that host models and datasets have gained widespread popularity.
-For example, Figure 1 shows the number of models and datasets (in a million scales) on Hugging Face, a popular AI model hosting platform (Face, [n.
-d.]), starting from July 31st, 2024, to Jun.
-30th, 2025.
+For example, Figure 1 shows the number of models and datasets (in a million scales) on Hugging Face, the largest public AI model hosting platform (Face, [n.
+d.]), starting from July 2024 to June 2025.
 By the end of June 2025, it has reached over 1.8M models and 450K datasets.
@@
 Taking Hugging Face as an example, there are four types of such models.
-First, the fine-tuned models adapt base models for specific tasks by training on additional task-specific datasets (Zou et al., 2023).
-Second, the adapters models add lightweight and modular layers to the pre-trained models for specific tasks (Hu et al., 2023).
-Third, the quantization models trade off the precision in numerical computations for accelerating inference and reducing memory consumption (e.g., using less precise model parameters) (Velingker et al., [n.
+First, fine-tuned models adapt base models for specific tasks by training on additional task-specific datasets (Zou et al., 2023).
+Second, adapters models add lightweight and modular layers to the pre-trained models for specific tasks (Hu et al., 2023).
+Third, quantization models trade off precision in numerical computations for accelerating inference and reducing memory consumption (e.g., using less precise model parameters) (Velingker et al., [n.
 d.]).
-Fourth, the merge models integrate multiple specialized or fine‑tuned parameter sets into a single unified model by combining weights or configurations (e.g., via weighted averaging or parameter stitching), enabling support for multiple tasks or domains without separate deployments (Akiba et al., 2025).
+Fourth, merged models integrate multiple models into a single unified model by combining weights or configurations enabling support for multiple tasks or domains without separate deployments (Akiba et al., 2025).
 Besides models, such platforms also host many datasets used for training and fine-tuning the previously discussed models (Röttger et al., 2024).
-Figure 1. The number of models and datasets (in millions) on Hugging Face from July 31, 2024, to June 30, 2025.
 1.1. Motivation
-As existing LLMs are often built from base models or other pre-trained models and using external datasets, they can inevitably inherit vulnerabilities, biases, or malicious components that exist in previous models or datasets.
-Therefore, understanding these components’ origin and development process can help better detect potential risks, improve model fairness, and ensure compliance with regulatory frameworks.
-Motivated by that, this project aims to study such relationships between models and datasets.
+As existing LLMs are often built from base models or other pre-trained models and using external datasets, they can inevitably inherit vulnerabilities, biases, or malicious components from previous models or datasets.
+Thus, understanding these components’ origin and provenance can help better detect potential risks, improve model fairness, and ensure compliance with regulatory frameworks.
+Figure 1. The number of AI models and datasets (in a million scale) on Hugging Face from July 2024 to June 2025.
+Motivated by that, this paper aims to study such relationships between models and datasets.
 They are the central parts of the LLM supply chain (Wang et al., 2024), which refers to the entire lifecycle of developing, training, and deploying LLMs, similar to a traditional supply chain in manufacturing or software development (Čolaković et al., 2021; Anonymous, 2022; Chou et al., 2006; Sonatype, 2015).
-The LLM supply chain can help to identify critical insights for both model evolution and dataset origin, as discussed below.
+Such a supply chain can help to identify critical insights for both model evolution and dataset origin, as discussed below.
 Model evolution.
-The study of the supply chain of LLMs gives a clear overview of how LLMs evolve from base models to fine-tuned variants, adapter integration, quantization models, and merged models that combine multiple specialized parameter sets into a single release.
+The study of the LLM supply chain gives a clear overview of how LLMs evolve from base models to fine-tuned variants, adapter integration, and quantization models.
 With that, one can easily keep track of them.
-For example, a use case is when a security or bias risk is found in one LLM, and we can quickly locate the potential models that might have the same problems.
+For example, a use case is when a security vulnerability is found in one LLM, and we can quickly locate the potential models that might have the same issues.
 Dataset origin.
-This supply chain can help to understand the datasets’ origins used for training different models (Shen et al., 2023).
+This supply chain can also help to understand the datasets’ origins used for training different models (Shen et al., 2023).
 Dataset origin refers to the source from which the data is collected.
 For example, for a fine-tuned model, we not only care about which dataset is used for fine-tuning but also what other datasets are involved in training the previous model.
-Understanding such dataset origin helps to ensure that the dataset used is reliable, ethical, legally compliant, and relevant to the task.
+Understanding such dataset origin helps to ensure that the dataset used is reliable, and legally compliant.
 1.2. Contribution
-Our contributions are mainly threefold.
+Our main contributions are threefold.
 First, we design a methodology to systematically collect the supply chain information of LLMs.
-In this paper, we mainly study the most popular ML platform, i.e., Hugging Face, but the same strategy also applies to other LLM platforms.
-In particular, we use the APIs from the ML platform to collect the metadata about the hosted model and dataset.
-To that end, we collected a large dataset, covering all the models (1.8M) and datasets (450K) as of June 30, 2025.
-Second, with the collected metadata, we designed a new graph, named LLM supply chain graph, to model the relationships between models and datasets.
-It is a directed heterogeneous graph where a node denotes different types of datasets and models (including base, fine-tune, adapter, quantization, and merge models).
-An edge denotes the dependency relationship between them, including dataset-dataset, and model-dataset relationships.
+In this paper, we mainly study the most popular AI platform, i.e., Hugging Face, but the same strategy applies to other platforms.
+In particular, we use the APIs from the AI platform to collect the metadata about the hosted model and dataset.
+To this end, we collected a large dataset as of June 30, 2025.
+Second, with the collected metadata, we created a new graph, named LLM supply chain graph, to model the relationships between models and datasets.
+It is a directed heterogeneous graph where a node denotes different types of datasets and models (including base, fine-tune, adapter, quantization, and merge).
+An edge denotes the dependency relationship between them, including model-model, dataset-dataset, and model-dataset relationships.
 Together, this graph is able to accurately capture the LLM supply chain information.
-To this end, we constructed a large graph with 397,376 nodes and 453,469 edges.
-An anonymized version of the complete graph is publicly available at GitHub [^1] Anonymized LLM supply chain graph: https://github.com/huggingface00/HuggingGraph.
-Third, with this graph, we perform different types of analysis, including model-dataset relationship analysis.
-We answer five research questions, including (i) the properties of the LLM supply chain graph, (ii) the structural analysis, (iii) the supply chain relationships between datasets, (iv) the supply chain relationships between the models and datasets, and (v) $\Delta$-based temporal update and tracking.
+To this end, we constructed a large graph with 402,654 nodes and 462,524 edges.
+An anonymized version of the complete graph is publicly available at GitHub [^1] LLM supply chain graph: https://github.com/huggingface00/HuggingGraph.
+Third, with this graph, we perform different types of analysis, including forward and backward analysis.
+We study seven research questions, including (i) the properties of the LLM supply chain graph, (ii) structural analysis, (iii) supply chain relationships between AI models, (iv) supply chain relationships between datasets, (v) supply chain relationships between models and datasets, (vi) dynamic update evaluation, and (vii) extension to other AI platforms.
 2. Preliminary
@@
 In addition, LLMs can undergo adaptation, quantization, and fine-tuning, a process where they are tuned with domain-specific datasets to maximize performance on specific tasks (VM et al., 2024), thus improving their accuracy and applicability.
-Note that this study mainly focuses on the relationships between various types of models and datasets, which are the central parts of the whole LLM supply chain ecosystem.
-We hope this study can not only provide insights on the relationships between LLM models and datasets but also raise awareness and future research interests in the LLM supply chain ecosystem.
+This study mainly focuses on the relationships between models and datasets, which are the central parts of the whole LLM supply chain ecosystem.
+We hope this study can not only provide insights on the LLM supply chain but also raise awareness and future research interests in this direction.
 3. Methodology
-This section discusses the main methodologies, including data collection, LLM supply chain graph construction, and analysis.
-Table 1. The APIs used to extract data from Hugging Face.
-API Name Description
-Model hub API Access the model hub to list, search, and download models and metadata.
-Dataset API Access the datasets for discovery, metadata retrieval, and downloading.
-Metrics API Access metrics for model evaluation, e.g., metric discovery, metadata retrieval, and calculation.
-Search API Search the name, tag, or other metadata for models and datasets.
 3.1. LLM Supply Chain Information Collection
@@
 Fortunately, platforms like Hugging Face provide some APIs that allow us to access the model and dataset and collect their metadata information, which can be used to construct the supply chain ecosystem.
-Table 1 summarizes four types of APIs we used.
-In particular, (i) the model hub APIs allow one to access the hub of existing models, including searching and downloading the model and its metadata.
-(ii) The dataset APIs allow one to access the datasets for discovery, metadata retrieval, and downloading.
-(iii) The metrics APIs allow one to access the metrics for model evaluation, including metric discovery, metadata retrieval, and calculation.
-(iv) In addition, one can use search APIs to search the name, tag, or other metadata for models and datasets.
-With the help of these APIs, we are able to collect the metadata of the hosted models and datasets.
-In particular, we collected a large dataset, covering the models (330K) and datasets (66K) as of June 30, 2025.
+Table 1 summarizes the four types of APIs we used.
+In particular, (i) the model hub APIs allow access to the hub of existing models, including searching and downloading the model and its metadata.
+(ii) The dataset APIs allow access to the datasets for discovery, metadata retrieval, and downloading.
+(iii) The metrics APIs allow access to the metrics for model evaluation, including metric discovery, metadata retrieval, and calculation.
+(iv) One can use the search APIs to search the name, tag, or other metadata for models and datasets.
+Table 1. The APIs used to extract data from Hugging Face.
+API Name Description
+Model hub Access the model hub to list, search, and download models and metadata.
+Dataset Access the datasets for discovery, metadata retrieval, and downloading.
+Metrics Access metrics for model evaluation, e.g., metric discovery, metadata retrieval, and calculation.
+Search Search name, tag, or other metadata for model and dataset.
 Handling missing information.
 Accurate construction of the LLM supply chain graph critically depends on the quality of metadata from the LLM platforms (e.g., Hugging Face), which might suffer from missing or incomplete data.
-To address this limitation, we design the following two techniques, i.e., cross-reference links and textual pattern extraction.
+To address this limitation, we apply the following two techniques, i.e., cross-reference links, and textual pattern extraction.
 (i) Cross-reference links.
@@
 This scraping step complements API-based extraction and is only employed when reverse dependency data is otherwise inaccessible.
-Figure 2. An example subgraph centering on the base model “Meta-llama” from the complete LLM supply chain graph.
 (ii) Textual pattern extraction.
@@
 For example, the textual phrase like “fine-tuned from Llama-2” contains the fine-tuned keyword, which implies from which model this model is actually fine-tuned.
-Similarly, we also look into other words, such as “train”, and “adapt”.
+Similarly, we also look into other words, such as, “train”, and “adapt”.
 3.2. LLM Supply Chain Graph
 With the collected metadata, we construct a directed heterogeneous graph to model the LLM supply chain.
-In this graph, a node denotes different types of datasets and models, including base, fine-tuned, adapter, and quantization models.
+In this graph, a node denotes different types of datasets and models, including base, fine-tune, adapter, quantization, and merge models.
 An edge denotes the dependency relationship between them, including model-model, dataset-dataset, and model-dataset relationships.
-Figure 2 shows a simplified supply chain subgraph centering on the base model Meta-llama.
-(i) Dataset-dataset relationship.
-The datasets within the LLM supply chain might overlap, build upon, or extend from each other–e.g., a dataset may be a subset or modified version of another.
+Figure 2 shows a simplified supply chain subgraph centering on a base model Meta-llama.
+(i) Model-model relationship.
+To further identify the supply chain relationship, we will check the relevant data fields.
+In particular, given a model in Hugging Face, there are data fields “finetune” that show which models are fine-tuned from this model.
+Similarly, “adapter”, “quantization”, and “merge show which models are adapted, quantized or merged from them, respectively.
+With such information, we can construct the supply chain relationship between the models.
+As shown in Figure 2, model Llama-3.3-70B is fine-tuned from the base model Meta-llama.
+Then, it is used by the models Doctor-Shotgun and Llama-3.3-70B-4bit-Vision to generate an adapter and quantization model, respectively.
+(ii) Dataset–dataset relationship.
+The datasets within the LLM supply chain might overlap, build upon, or extend from each other.
+For example, a dataset may be a subset or modified version of another.
 To capture such information, we connect them with two types of edges.
-(1) Subset relationship.
-These arise when a dataset is explicitly documented as a subsample or partition of another.
+(1) Subset relationships arise when a dataset is explicitly documented as a subsample or partition of another.
 For instance, a dataset named “C4_200M” is described as a subset of “C4”.
-(2) Modified version.
-These represent updates or enhanced variants of existing datasets.
-For example, “TruthfulQA_v2” incorporates factual corrections and improvements over an earlier version, “TruthfulQA_v1”.
-(ii) Model-dataset relationship.
-To capture the model and dataset dependency, we use the metadata information from both models and datasets.
-From the metadata of a model, it might include the datasets used to train it.
+(2) Modified versions represent updates or enhanced variants of existing datasets.
+For example, “TruthfulQA_v2” incorporates corrections and improvements over an earlier version, “TruthfulQA_v1”.
+(iii) Model-dataset relationship.
+To capture the model and dataset dependency, we use the metadata from both models and datasets.
+The metadata of a model might specify the datasets used for training or adapting.
 However, not all the models disclose such information.
-To find more information, we observe the metadata of a dataset, including a data field on “trained_fine_tune_models” on this dataset.
-Thanks to that, we are able to capture the accurate model and dataset relationship.
-In Figure 2, the datasets The Pile and Awesome-Chatgpt-prompt have directed edges to model Meta-Llama, meaning that both datasets are used to train the model.
+To capture more information, we find the metadata of a dataset contains a data field of “trained_fine_tune_models” on this dataset.
+Thanks to that, we are able to capture the accurate model and dataset relationships.
+In Figure 2, the datasets The Pile and Chatgpt-prompt have directed edges to model Meta-Llama, meaning that both datasets are used to train the model.
+Figure 2. An example subgraph centering on base model “Meta-llama” from the complete LLM supply chain graph.
 3.3. Supply Chain Graph Analysis
 This supply chain graph can help to understand the transformational processes of the models and datasets.
-In particular, we can understand how base models evolve into their variants, including fine-tuned, adaptive, and quantized models, and vice versa.
+In particular, we can understand how base models evolve into their variants, including fine-tuned, adaptive, quantized, or merged models, and vice versa.
 Similar observations can be made for datasets.
@@
 In particular, we mainly perform two types of analysis, i.e., forward and backward analysis.
+Forward analysis is the method of traversing the supply chain graph following the dependency edges of a chosen node in a forward-going way.
+This node (known as the root/source node) can be a dataset, a base model, a fine-tuned model, an adapter, or a quantized model.
+In particular, given a source node, we apply the graph traversal algorithm (e.g., breadth-first search (BFS) (Tarjan, 1972)) to traverse all the nodes (including both models and datasets) in a level-by-level pattern.
+To that end, this forward analysis will identify all the nodes that are reachable from the source node.
+Model analysis example.
+In Figure 2, to analyze the forward supply chain of Meta-Llama, we trace its development from training datasets to subsequent adaptations.
+This model is fine-tuned into Llama-3.3-70B, enhancing its capabilities for specific tasks.
+Summarizing the forward supply chain of Meta-Llama, we identify the following four distinct forward supply chains: (i) Base model (Meta-llama) $\rightarrow$ fine-tuned model (Llama-3.3-70B) $\rightarrow$ another fine-tuned model (Llama3.3-70B-Vision).
+(ii) Base model (Meta-llama) $\rightarrow$ adapted model (Doctor-Shotgun).
+(iii) Base model (Meta-llama) $\rightarrow$ fine-tuned model (Llama-3.3-70B) $\rightarrow$ quantization model (Llama-3.3-70B-4bit).
+(iv) Base model (Meta-llama)) $\rightarrow$ merged model (MistLlama).
+These pathways illustrate the evolution trajectory of the base model Meta-Llama showcasing its progressive specialization and adaptation for various tasks.
 Dataset analysis example.
-For the dataset, our supply chain analysis shows how different datasets connect and form a new dataset and work together.
-This combination creates flexible resources that improve how models perform in various areas.
+For the dataset, our supply chain analysis shows how different datasets connect and form a new dataset.
+This combination creates flexible resources that show how models perform in various areas.
 In Figure 2, the dataset The Pile is composed of multiple subsets, including Wikimedia, Arxiv, Openwebtext2, and Pubmed Central.
 Together, these datasets form a unified corpus that serves as training data for models like Meta-LlaMA.
+Backward analysis, on the other hand, is the method of traversing the supply chain graph following the edges in a backward way.
+We accomplish this by traversing the directed graph also with BFS (Knuth, 1974) in reverse, starting from the selected node and following the incoming edges.
+To that end, this backward analysis will identify all the nodes that can reach the source node.
 Model and dataset analysis example.
-In Figure 2, to analyze the backward supply chain of the model RBot70Bv4, we trace its lineage back through its development stages.
-This model is fine-tuned from Unsloth, which in turn originates from its base model, Meta-Llama, and the datasets used to train the base model are The Pile and Awesome-Chatgpt-prompts.
-Through this analysis, we establish the backward path, starting from the target model—RBot70Bv4—and tracing back to its base model, Meta-Llama, revealing the dependencies and transformations involved in its development.
-Similarly, we can analyze the datasets.
-3.4. Temporal Update and Tracking
-To ensure HuggingGraph remains an accurate and evolving representation of the LLM ecosystem, we have implemented a (i) $\Delta$-based batch update mechanism and (ii) a versioned archival system.
-$\Delta$-based graph update.
-To ensure HuggingGraph reflects the evolving nature of the LLM ecosystem, we implement a $\Delta$-based batch update mechanism.
-Instead of reconstructing the graph from scratch, we compute incremental updates by comparing successive versions.
-Let $G_{t}$ denote the graph at time $t$, and define the $\Delta$ as $\Delta_{t}=G_{t}-G_{t-1}$, where $\Delta_{t}$ captures all added or removed nodes and edges.
-This approach offers both computational scalability and storage efficiency.
-At scheduled intervals (e.g., daily), we collect metadata from Hugging Face and compute the differences.
-For instance, between June 25–July 15, 2025, the platform added 104,380 new models ($\sim$3,866/day), 20,034 new datasets ($\sim$742/day), and approximately 854 dependency edges/day.
-Versioning and longitudinal tracking.
-To enable temporal and historical analyses, each update generates a versioned snapshot of the graph, archived with a timestamp and associated change log.
-This log captures additions, deletions, and dependency modifications, such as new edges from TruthfulQA v2 to TruthfulQA v1.
-This ensures that HuggingGraph remains a living resource that adapts to ongoing developments in LLM research and deployment.
-4. Experiment and Finding
-To deeply understand the relationships between models and datasets, we aim to answer seven critical research questions (RQs).
-We believe they can offer valuable insights into the supply chain of the LLM ecosystem and potentially pave the way for future LLM development and deployment.
-  • RQ #1: What are the properties of the LLM supply chain graph?
-  • RQ #2: What structural patterns emerge from the LLM ecosystem?
-  • RQ #3: What are the supply chain relationships between datasets?
-  • RQ #4: What are the relationships between models and datasets?
-  • RQ #5: What insights can be gained as the graph temporally updating?
+In Figure 2, analyzing the backward supply chain of model RBot70Bv4, we trace its lineage through its development stages.
+This model is fine-tuned from Unsloth, which in turn originates from its base model, Meta-Llama, and the datasets used to train the base model are The Pile and Chatgpt-prompts.
+Through this analysis, we establish the backward path, starting from the target model, i.e., RBot70Bv4, and tracing to its base model, Meta-Llama, revealing dependencies and transformations involved in its development.
+Similarly, we can analyze datasets.
+3.4. Accommodating Dynamic Update
+The hosted models and datasets on AI platforms are growing fast as new models are being developed every day.
+For example, between June 25 and July 15, 2025, we observed 80,703 new models (approximately 3,843 per day), 27,405 new datasets (approximately 1,305 per day) on Hugging Face only, which is just one of the many AI platforms.
+Therefore, we need to accommodate the dynamic update to accurately manage and analyze the AI supply chain.
+Particularly, HuggingGraph accommodates the dynamic update in three steps.
+(i) Scoping updated models or datasets.
+At a time $t$, we keep a copy of the hosted models and datasets with their IDs.
+When we evolved to $t+1$, we get another copy of the hosted models and datasets with their IDs.
+The difference between them shows the updated models and datasets, including newly added or deleted.
+In our current implementation, we are keeping the update on a daily basis.
+(ii) Metadata collection for the updated models or datasets.
+For the identified updated models or datasets, we will collect their metadata using the same strategy as discussed in Section 3.1.
+To this end, we get the updated dependencies between models and datasets.
+That is, for the update at time $t+1$ compared to $t$, it can be represented as $\Delta_{t+1}$.
+(iii) $\Delta$-based dynamic graph update.
+Given the newly updated dependency $\Delta_{t+1}$, and let $G_{t}$, $G_{t+1}$ denote the graph at time $t$, $t+1$, respectively, then $G_{t+1}=G_{t}\cup\Delta_{t+1}$.
+4. Experiments and Finding
+To deeply understand the relationships between models and datasets, we study seven critical research questions (RQs), which can offer valuable insights into the supply chain of the LLM ecosystem and potentially pave the way for future LLM development.
+  • RQ #1: What are the properties of LLM supply chain graph?
+  • RQ #2: What structural patterns emerge?
+  • RQ #3: What are the supply chain between LLM models?
+  • RQ #4: What are the supply chain between datasets?
+  • RQ #5: What are the relationships between models and datasets?
+  • RQ #6: What insights can be gained from the dynamic updates?
+  • RQ #7: How can HuggingGraph be applied to other platforms?
 4.1. RQ #1: Supply Chain Graph Properties
-Figure 3.
-(a) Indegree distribution and (b) Outdegree distribution.
-The X-axis represents the number of indegrees and outdegrees, respectively, and the Y-axis represents the number of vertices in a logarithmic scale.
-This research question aims to understand the critical properties of the LLM supply chain graph, i.e., graph basics and degree distribution.
+Figure 3. (a) Indegree distribution and (b) Outdegree distribution. X-axis represents degrees, and the Y-axis represents the number of vertices in a logarithmic scale.
+This research question aims to understand the critical properties of LLM supply chain graph, i.e., graph basics and degree distribution.
 Graph basics.
-The collected supply chain graph is a large directed heterogeneous graph with 397,376 nodes and 453,469 edges as of June 30th, 2025.
-In particular, there are six different types of nodes, including 25,717 base models, 115,103 fine-tuned, 79,240 adapters, 98,125 quantization models, 13,027 merges and 66,164 datasets.
-The average degree is about 1.14, denoting that it is a very sparse graph.
-The degree distribution of a graph describes how node degrees (the number of edges connected to a node) are distributed across the graph.
-Figure 3(a) and (b) illustrate the indegree and outdegree distribution of the graph, respectively.
-We show not only the total distribution but also the distribution of five types of nodes, including base models, fine-tuned models, adapter models, quantization models, and datasets.
+The collected supply chain graph is a medium-scale directed heterogeneous graph with 402,654 nodes and 462,524 edges as of June 30th, 2025.
+In particular, there are six different types of nodes, including 28,384 base models, 115,211 fine-tuned, 79,254 adapters, 98,143 quantization models, 13,028 merges, and 68,634 datasets.
+The average degree is about 1.15, denoting that it is a very sparse graph.
+Furthermore, we identified substantial metadata missing.
+As of June 30, 2025, among 1.8 million models, only 50,156 (2.76%) include a model tree, while $\sim$ 550K models lack any metadata beyond their name, with nearly 400K entries empty or invalid.
+Similarly, of the 450K datasets, only 68,634 (15.26%) provide datacards, leaving $\sim$ 380K datasets without any usable metadata.
+This highlights a broader issue in the AI community, where a significant number of models and datasets lack consistent and structured documentation on the supply chain.
+This reflects the need of more transparent disclosure.
+Degree distribution of a graph describes how node degrees (the number of edges connected to a node) are distributed across the graph.
+Figure 3 illustrates the indegree and outdegree distribution of the graph.
+We show not only the total distribution but also the distribution of six types of nodes, including base models, fine-tuned models, adapter models, quantization models, merged models, and datasets.
 We made two interesting observations.
@@
 The outdegree distribution follows a similar pattern but may differ in specific cases (e.g., adapters seem to have a more restricted degree distribution).
-The power-law behavior suggests that most nodes have low degrees, while a few central nodes (hubs) dominate the graph.
-In particular, the model images from macrocosm-os has the highest indegree value of 550, and Mistral-7B-v0.1 from mistralai has the highest outdegree value of 1,093.
+The heavy-tailed behavior suggests that most nodes have low degrees, while a few central nodes (hubs) dominate the graph.
+In particular, the dataset macrocosm-os/images has the highest indegree value of 550, and Mistral-7B-v0.1 from “mistral AI” has the highest outdegree value of 1,093.
 Specifically, the base models act as high-degree hub nodes as they are heavily used by other task-specific models.
-To this end, we can conclude that this graph is a power-law graph, also known as a scale-free graph.
-That is, a graph whose degree distribution follows a power law distribution.
-In other words, a small number of nodes have many connections, while most nodes have a few connections.
-(ii) We observe that all the base models have at least one outdegree as they are used by the task-specific models.
-For the dataset nodes, there are 20,093 datasets with zero indegrees, meaning they are the minimum datasets that can no longer be partitioned.
-Finding #1: The LLM supply chain graph is a large, sparse graph that exhibits a power-law degree distribution, where a small number of high-degree base models and datasets serve as central hubs.
+Finding #1: The LLM supply chain graph is medium-scale, sparse, and heavy-tailed distribution.
+However, a significant number of models and datasets lack metadata, highlighting the need for more transparent supply chain documentation.
 4.2. RQ #2: Supply Chain Structural Analysis
+Figure 4. Cumulative distribution function (CDF) of WCC size.
 This research question aims to understand the topology and evolution of the LLM supply chain. To achieve that, we analyze the structural properties with connectivity and community analysis.
-Connectivity analysis in a graph is crucial for understanding how well different nodes are linked and helps identify critical nodes and edges that maintain overall connectivity.
-This graph is a directed acyclic graph (DAG) because there are no dependencies in a cycle form, which also means there are no strongly connected components.
-Because of that, we only perform weakly connected components (WCC) analysis, and the total number of WCCs in our supply chain graph is 44,417.
-Figure 4. Cumulative distribution function (CDF) of WCC size.
+Connectivity analysis analyzes the structural connectivity information.
+We performed weakly connected components (WCC), which identifies maximal sets of nodes that remain connected when edge directions are ignored.
+The total number of WCCs in our supply chain graph is 44,908.
 Figure 4 shows the cumulative distribution function (CDF) of the WCC distribution.
 We made two interesting observations.
-(i) The largest WCC covers 244,063 nodes, accounting for 61.4% of all the nodes.
+(i) The largest WCC covers 247,244 nodes, accounting for 61.8% of all the nodes.
 It reflects the dense interconnections that pervade the ecosystem.
@@
 In the largest WCC, major models are included, such as Gemma-2B, DistilBERT, and GPT-2.
-(ii) In contrast, the remaining 153,313 WCCs collectively hold 31.4% of the nodes, with most having 1, 2, to 3 nodes.
-This indicates a fragmented outer edge characterized by specialized models, less frequent datasets, or active experimental projects.
+(ii) In contrast, the remaining WCCs collectively hold 38.6% of the nodes (i,e,.155,410 nodes), with most having 1, 2 or 3 nodes.
+This indicates a fragmented outer edge characterized by specialized models, rare datasets, or active experimental projects.
 The prevalence of these small, isolated pieces suggests niche attempts that lack integration with the overall system.
-For example, tblard/tf-allocine (French sentiment), distilbert-base-uncased-finetuned-sst-2-english (sentiment analysis), and PolyAI/minds14 (banking queries) remain disconnected due to limited reuse or insufficient metadata.
-Community analysis is the process of identifying groups or clusters of nodes within a graph that are more densely connected internally than with the rest of the graph.
+For example, zhongqy/RMCBench (benchmarking dataset) and yigit69/bert-base-uncased-finetuned-rte-run_3 (recognizing textual entailment task) remain disconnected due to limited reuse or insufficient metadata.
+We also computed strongly connected component (SCC), a maximal subgraph in which every node is reachable from every other via directed edges, identifying 398,198 SCCs.
+Remarkably, only 591 of these are non‑trivial (size ¿ 1), collectively encompassing 2,169 nodes (0.54% of the graph).
+The largest SCC comprises 478 dataset nodes, among them tree‑of‑knowledge and OpenHermes‑2.5, forming a tightly‑knit cluster.
+By contrast, the remaining 99.46% of nodes each reside in trivial (size‑1) SCCs.
+Community detection is the process of identifying groups or clusters of nodes within a graph that are more densely connected internally than with the rest of the graph.
 In the context of the LLM supply chain, it reveals semantically or functionally aligned subgraphs that reflect the patterns of reuse and task-specific models or datasets.
@@
 Modularity is a measure of how well a graph is divided into communities, where a high modularity score indicates that most edges fall within communities rather than between them.
-It compares the actual density of edges inside communities to the expected density if edges were placed at random.
 Table 2 summarizes the top-10 communities detected using the Louvain method.
-Our analysis reveals an exceptionally modular structure, with each of the top communities achieving a high modularity score of 0.97, indicative of strong intra-community connectivity.
+Our analysis reveals an exceptionally modular structure, with each of the top communities achieving a high modularity score of 0.96, indicative of strong intra-community connectivity.
 Collectively, these communities span a wide range of functional domains, underscoring the presence of well-defined, task-aligned clusters within the ecosystem.
 We made two interesting observations.
-(i) The largest community (ID 1) consists of 9,158 nodes and attains a modularity score of 0.97, indicating extremely cohesive internal structure.
-It revolves around base models such as OLMoE and CausalLM, and general-purpose datasets like prompt-perfect.
-This suggests a densely connected nucleus facilitating widespread reuse and fine-tuning—effectively functioning as the backbone of the LLM ecosystem.
+(i) The largest community consists of 9,390 nodes and attains a modularity score of 0.96, indicating an extremely cohesive internal structure.
+It revolves around base models like OLMoE and CausalLM, and general-purpose datasets like prompt-perfect.
+This suggests a densely connected cluster facilitating widespread reuse and fine-tuning.
 (ii) Several other communities reflect clear task-based segmentation.
-For instance, Community 3 ( 7,058 nodes, modularity 0.97) focuses on benchmarking, with models like Wanxiang and datasets such as marco-o1.
-Similarly, Community 6 (5,030 nodes, modularity 0.97) centers on instruction-tuning, connecting models like MedLlama-3-8B with curated datasets like dpo-mix.
-Table 2. Top-10 Louvain communities in the HuggingGraph.
+For instance, Community 3 (7,388 nodes, modularity 0.96) focuses on benchmarking, with models like Wanxiang and datasets like marco-o1.
+Similarly, ID 5 (4,554 nodes, modularity 0.96) centers on instruction-tuning, connecting models like MedLlama-3-8B with curated datasets like dpo-mix.
+Table 2. Top-10 Louvain communities sorted by size.
 ID Size E.g. models E.g. datasets Modularity
-1 9158 OLMoE, CausalLM prompt-perfect 0.97
-2 7058 qwen2.5_math Marco-o1 0.97
-3 6987 Wanxiang smartllama3.1 0.97
-4 5696 tinyllama Llama-1B 0.97
-5 5355 Qwen2.5-32B Matter-0.2 0.97
-6 5030 MedLlama-3-8B dpo-mix 0.97
-7 4297 Electra, ArliAI MixEval 0.97
-8 4133 aesqwen1.5b llava 0.97
-9 4067 bert TORGO 0.97
-10 3932 Mistral vicuna_format 0.97
-Finding #2: The LLM supply chain graph reveals a structure with a densely connected core and a fragmented periphery.
-A single dominant component encompasses the majority of nodes, supporting efficient model reuse and dependency tracing, while many smaller, isolated components reflect specialized or under-integrated efforts.
-Community analysis shows strong modularity, with nodes clustering around shared tasks and functional roles.
-4.3. RQ #3: Dataset Supply Chain Analysis
+1 9,390 OLMoE, CausalLM prompt-perfect 0.96
+2 7,388 qwen2.5_math Marco-o1 0.96
+3 6,989 Wanxiang smartllama3.1 0.96
+4 6,813 tinyllama Llama-1B 0.96
+5 5,163 Qwen2.5-32B Matter-0.2 0.96
+6 4,554 MedLlama-3-8B dpo-mix 0.96
+7 4,262 Electra, ArliAI MixEval 0.96
+8 3,947 aesqwen1.5b llava 0.96
+9 3,829 bert TORGO 0.96
+10 3,828 Mistral vicuna_format 0.96
+Finding #2: The LLM supply chain graph features a dominant core (61.8% of nodes), while high modularity (0.96) reveals task-aligned, semantically coherent communities amid a fragmented periphery.
+4.3. RQ #3: Supply Chain Analysis of LLM Models
+This research question aims to provide a holistic view of the dependencies between the models within the LLM supply chain, particularly from both base and task-specific models.
+Base model impact.
+We would like to understand the impact of base models.
+Here, we quantify the impact of a base model as the number of task-specific models that depend on it.
+The more dependencies, the larger the impact it has.
+We start with a base model and perform forward analysis by computing a breadth-first search (BFS) following the outgoing edges to get that from our LLM supply chain graph.
+This leads to a forward subgraph, which denotes all the models that depend on the base model, including fine-tuned, adapted, quantized, or merged models.
+Table 3 shows the top-10 base models sorted by the forward subgraph size, which is the number of impacted task-specific models.
+We make two interesting observations.
+(i) A base model can significantly impact the LLM supply chain ecosystem.
+For example, Llama-3.1-8B is a base model from Meta used for efficient text generation, code assistance, and research (He et al., 2024).
+Due to its relatively small size, which allows for deployment in resource-constrained environments, making advanced AI accessible to broader stakeholders (Grattafiori et al., 2024).
+It has generated up to 7,544 models, including 1,710 fine-tuned versions, 1,542 adapters, 3,473 quantizations, and 1,693 merged models tailored to specific tasks.
+(ii) In terms of fine-tuning, the base model Mistral-7B-v0.1 has been fine-tuned the most times, totaling 2,105.
+It is a faster and lighter version of the Mistral model from Mistral AI.
+It is obtained by training on a large corpus with grouped-query and sliding window attention by Mistral AI, delivering efficient text generation, NLP, and code assistance on consumer hardware for low-latency applications like chatbots and text classification (Thakkar and Manimaran, 2023).
+Table 3. Top-10 base models sorted by forward subgraph size.
+Base model Total
+Fine-
+tune
+Adapter
+Quanti-
+zation
+Merge Level
+Llama-3.1-8B 7544 1710 1542 3473 1693 25
+Mistral-7B-v0.1 6744 2105 2187 1435 1254 27
+Qwen2.5-7B 6733 1972 1764 2516 1132 11
+Meta-Llama-3-8B 5633 967 1511 2220 1967 21
+Llama-3.1-70B 4063 698 281 2075 2519 11
+Qwen2.5-32B 3909 1086 158 2311 1049 12
+Qwen2.5-1.5B 3645 1300 1290 949 248 8
+Qwen2.5-0.5B 3521 1669 1006 810 46 11
+Qwen2.5-14B 3362 726 411 1880 1166 15
+Meta-Llama-3-8B-Instruct 3118 640 405 1394 1305 34
+Task-specific model analysis.
+Given a task-specific model, we want to understand how it evolves, meaning what other models it has relied on.
+To achieve that, for each task-specific model, we perform a backward analysis by running a BFS following the incoming edges.
+To that end, the derived subgraph shows the models it relies on.
+Table 4 shows the top 10 task-specific models sorted by the backward subgraph size, which is the number of models they rely on.
+We make two interesting observations.
+(i) A fine-tuned model, command-r-1-layer, illustrates the depth and complexity of transformations in the LLM supply chain.
+This model operates in bfloat16 (BF16) precision for efficient text generation and natural language understanding (Pignatelli et al., 2024), originates from the base model cc4ai, and has undergone extensive lineage evolution before reaching its final form.
+Specifically, it depends on 40 upstream artifacts, including 39 other fine-tuned models, and spans 39 transformation levels in its backward lineage chain, as detailed in Table 4.
+(ii) We observed that adapters are mainly used for lightweight fine-tuning and merges for model integration, but task-specific models like command-r-1-layer, as optimized standalone derivatives, do not evolve from adapters or merges in their backward lineage (Pignatelli et al., 2024).
+Finding #3: Base models like Llama-3.1-8B dominate the LLM supply chain, spawning thousands of derivatives, while task-specific models such as command-r-1-layer exhibit deep dependencies with other task-specific variants but avoid adapters or merges.
+4.4. RQ #4: Dataset Supply Chain Analysis
+Table 4. Top-10 models sorted by backward subgraph size.
+Model
+Model
+Type
+Total
+Fine-
+tune
+Quanti-
+zation
+Level Base Model
+command-r-1-layer Finetune 40 39 0 39 c4ai
+KoModernBERT Finetune 21 20 0 20 ModernBERT
+t5-small Finetune 21 20 0 20 t5-small
+clinical_260k Finetune 20 19 0 19 clinical_180K
+t5-small-finetuned Finetune 17 16 0 16 t5-small
+clinical_300k Finetune 16 15 0 15 clinical_180K
+clinical_259k Finetune 16 15 0 15 clinical_180K
+LeoPARD-0.8.1 Finetune 16 2 13 15 DeepSeek-R1
+LeoPARD-0.8.2-4bit Quantization 16 1 14 15 DeepSeek-R1
+LeoPARD-0.8.1-4bit Quantization 16 1 14 15 DeepSeek-R1
 This research question aims to provide comprehensive insights into the different usages of the datasets within the LLM supply chain graph.
@@
 On one hand, one dataset can be created by combining a variety of other datasets from disparate sources.
-On the other hand, one dataset can serve as a building block for new datasets, which enables continuous improvement in machine learning assets.
-In the following, we discuss the impacts on both training and benchmarking datasets.
-Training dataset impact.
-In the LLM supply chain graph, 66,164 datasets act as training datasets that are used for model training or generating new datasets.
+On the other hand, one dataset can serve as a building block for new datasets.
+In the LLM supply chain graph, 68,634 datasets act as training datasets that are used for model training or generating new datasets.
 Table 5 presents the top 10 datasets ranked by the highest number of included and derived datasets.
@@
 Sitting at the top is macrocosm-os/images (Kido, 1961) from macrocosm.
-It includes 550 sub-datasets and is a general-purpose image modification corpus used in advance multimodal AI research and large-scale vision-language model training.
-Moreover, bespokelabs/Bespoke (Elliot et al., 2024) from bespokelabs includes 215 sub-datasets, is a high-quality synthetic dataset designed to enhance multimodal AI research.
-(ii) The other observation is that a single dataset can be included in many larger datasets, showing the huge overlaps between the datasets and also the models trained with them.
-The right two columns of Table 3 show the top 10 datasets sorted by the maximum number of derived datasets they have been included.
-The leading contributor, HuggingFaceH4 from HuggingFace has been involved in 989 derived datasets.
-It is primarily used for natural language understanding and generation tasks, including chatbots, text generation, code completion, reasoning, and multilingual processing, enabling the training and enhancement of models specialized in advanced conversational and instruction-following capabilities (Dietrich and Hollstein, 2025).
-Similarly, the CodeFeedback dataset from m-a-p has facilitated the creation of 857 derived datasets, is a collection of high-quality code instruction queries designed to enhance the training of large language models (LLMs) for code generation, debugging, and explanation tasks, enabling improved performance in complex programming scenarios(Meyers et al., 2018).
-Table 3. Top-10 datasets sorted by the # of included and # of derived datasets.
-Dataset
-# of included
-dataset
-Dataset
-# of derived
-datasets
+It is included in 550 datasets and is a general-purpose image modification corpus used in advanced multimodal AI research and large-scale vision-language model training.
+Moreover, bespokelabs/Bespoke (Elliot et al., 2024) from bespokelabs is part of 215 datasets, is a high-quality synthetic dataset designed to enhance multimodal AI research.
+Table 5. Top-10 datasets by # of included & derived datasets.
+Dataset # of included Dataset # of derived
 macrocosm-os/images 550 HuggingFaceH4 989
 bespokelabs/Bespoke 215 CodeFeedback 857
-databricks/databricks 137 MADLAD 850
-Open-Orca/OpenOrca 119 Capybara 823
+databricks/databricks 139 MADLAD 850
+Open-Orca/OpenOrca 121 Capybara 823
 nguha/legalbench 117 Glot500 804
-OpenAssistant/oasst1 111 dolphin-coder 756
-LDJnr/Capybara 104 SlimOrca 726
-kvn420/Tenro_V4.1 101 orca_dpo_pairs 687
-teknium/OpenHermes 98 gutenberg 675
-Anthropic/hh-rlhf 96 samantha 658
-Finding #4: The datasets play critical roles in model training with 66,164 datasets are contributing to model development through inclusion and derivation, as seen in macrocosm-os/images and HuggingFaceH4.
-4.4. RQ #4: Supply Chain between LLM Models and Datasets
+OpenAssistant/oasst1 112 dolphin-coder 756
+LDJnr/Capybara 108 SlimOrca 726
+kvn420/Tenro_V4.1 105 orca_dpo_pairs 687
+teknium/OpenHermes 104 gutenberg 675
+Anthropic/hh-rlhf 98 samantha 658
+(ii) The other observation is that a single dataset can be derived to many datasets, showing the huge overlaps between the datasets and also the models trained with them.
+The right two columns of Table 5 show the top 10 datasets sorted by the maximum number of derived datasets they have been included in.
+The leading contributor, HuggingFaceH4 from HuggingFace has been derived to 989 datasets.
+It is primarily used for natural language understanding and generation tasks, including chatbots, text generation, code completion, reasoning, and multilingual processing, enabling the training and enhancement of models specialized in advanced conversational and instruction-following capabilities (Dietrich and Hollstein, 2025).
+Similarly, the CodeFeedback dataset from m-a-p has facilitated the creation of 857 datasets, which is a collection of high-quality code instruction queries designed to enhance the training of large language models (LLMs) for code generation, debugging, and explanation tasks, enabling improved performance in complex programming scenarios(Meyers et al., 2018).
+Finding #4: Datasets play critical roles in training with 68,634 datasets contributing to model development through inclusion and derivation, as seen in macrocosm-os/images and HuggingFaceH4.
+4.5. RQ #5: Supply Chain between Models and Datasets
 This research question aims to explore the interconnections between models and datasets within the supply chain graph.
-It provides valuable insights from dual perspectives, as discussed below, including one dataset versus multiple models and one model versus multiple datasets.
+It provides valuable insights from dual perspectives, including one dataset versus multiple models and one model versus multiple datasets.
 One dataset versus multiple models refers to the case when a single dataset is used to train multiple models.
-Table LABEL:tab:dataset_model_1 shows the top 10 datasets based on the number of models trained on them.
+Table 6 shows the top 10 datasets based on the number of models trained on them.
 In particular, Mistral-7B-v0.1 takes the leading position and is a widely adopted open-source dataset known for its strong performance in general-purpose language understanding and generation tasks.
@@
 One model versus multiple datasets refers to the case when an LLM model is trained with multiple datasets.
-Table LABEL:tab:dataset_model_2 shows the top 10 models ranked by the number of datasets used for training.
-DeBERTa-ST-AllLayers-v3.1, a fine-tuned variant of the DeBERTa architecture, takes the top position, having been trained on 116 different datasets.
+We observed that DeBERTa-ST-AllLayers-v3.1, a fine-tuned variant of the DeBERTa architecture, takes the top position, having been trained on 116 different datasets.
 Its adapter-based counterpart, DeBERTa-ST-AllLayers-v3.1bis, also leverages the same number of datasets via adapter-based training, emphasizing modular reuse across tasks.
-In addition, models like static-similarity-mrl-mul-v1 and static-similarity-mrl-multilingual are both fine-tuned on 108 datasets, indicating their role in multilingual and multi-task similarity-based retrieval applications.
-The ModernBERT-base-embed and Llama-3.2-3B-Instruct families show strong dataset diversity as well, with 88 and 87 datasets respectively, across fine-tuning and quantized variants (e.g., GGUF format).
-Interestingly, most of these models are fine-tuned—specifically, 8 out of the top 10—highlighting a trend where base models are transformed into task-specific variants through diverse training datasets.
-This pattern suggests that fine-tuning remains a dominant strategy for adapting base models to downstream tasks across heterogeneous data sources.
-Finding #5: There exists a strong interdependence between models and datasets, where widely adopted datasets like Mistral-7B-v0.1 and TinyLlama-1.1B-v1.0 serve as the base for training hundreds of models across different adaptation techniques, while models such as DeBERTa-ST-AllLayers-v3.1 leverage diverse datasets to enhance their versatility, highlighting the critical role of dataset-model interactions in advancing AI capabilities.
 Table 6. Top-10 datasets sorted by # of models trained.
-Dataset
- Total
-Fine-
-tune
-Adap-
-ter
-Quanti-
-zation
-Merges
+Dataset Total Fine-tune Adapter Quantization Merges
 Mistral-7B-v0.1 1093 300 300 193 300
@@
 Starling-LM-7B-alpha 210 10 165 18 17
-Table 5. Top-10 models sorted by the # of datasets used for training.
-Model
-Model Type
-# of dataset used-
-for training
-DeBERTa-ST-AllLayers-v3.1 Fine tune 116
-DeBERTa-ST-AllLayers-v3.1bis Adapters 116
-static-similarity-mrl-mul-v1 Fine tune 108
-static-similarity-mrl-multilingual Fine tune 108
-ModernBERT-base-embed Fine tune 88
-Llama-3.2-3B-Instruct Fine tune 87
-Llama-3.2-3B-Instruct-GGUF Quantization 87
-DavidLanz-3.2-3B-Instruct Fine tune 87
-static-retrieval-mrl-en-v1 Fine tune 79
-XLMRoBERTaM3-CustomPoolin Fine tune 72
-4.5. RQ #5: Temporal Update and Tracking
-This research question aims to focus on the value of tracking fine-grained, time-based changes in the LLM supply chain.
-Our $\Delta$-based update mechanism (discussed in Section 3.4), produces a daily snapshot of the HuggingGraph, capturing precisely how many nodes and edges are added each day, how that translates into net growth, and what the busiest single day looks like.
-To better understand the implications of this tracking, we adopt both qualitative and quantitative perspectives on the evolving LLM ecosystem.
-Fine-grained temporal insights into the LLM ecosystem.
-We would like to understand how the LLM supply chain evolves on a day-to-day basis.
-Hence, we adopt this fine-grained temporal qualitative approach that highlights fluctuations in model and dataset activity.
-This strategy allows us to uncover ecosystem-level dynamics such as release bursts, contributor behavior, and category-specific trends over time.
-Figure 5 visualizes the daily $\Delta$-based growth of five key node categories (base, fine-tuned, adapters, quantized variants models, and datasets) from June 25 to July 15, 2025.
-We observe that, (i) the fine-tuned models dominate the daily activity, averaging over 1,700 new entries per day, followed by consistent contributions from adapters ($\sim$750/day) and datasets ($\sim$700/day).
-Clear spikes on specific days—e.g., July 7 and July 9—coincide with major releases such as the Mistral-Fusion-v3 fine-tuning wave and dataset refreshes like HFTime2025-News.
-(ii) Furthermore, adapter uploads peaked at 887 on June 28, while quantized variants reached 290 on July 9, driven by releases such as QWin-GGUF-7B.
-These patterns demonstrate HuggingGraph’s ability to capture evolving supply chain dynamics at a fine-grained, time-sensitive resolution, revealing ecosystem bursts and contributor behaviors in real time.
-Quantifying daily incremental graph growth.
-To complement our qualitative insights with measurable evidence, we adopt a quantitative approach that captures the scale and efficiency of daily changes to the LLM supply chain.
-This allows us to evaluate the performance of our $\Delta$-based update mechanism, ensuring support for large-scale, time-sensitive graph maintenance in a computationally efficient and auditable manner.
-Table 6 provides a detailed breakdown of daily additions, net changes, and peak updates across node and edge types from Jun.
-25 to Jul.
-15, 2025.
-(i) On average, Hugging Face added 126 new fine-tuned models per day (peaking at 2,090), followed by adapters and datasets with 41 and 39 daily insertions, respectively.
-Quantized and merged models contributed 16 and 1 additions per day, respectively.
-In total, 2,526 fine-tuned models and 785 datasets were added during this period.
-(ii) Each day also saw approximately 23 $\pm$ 6 new dependency edges—linking models to their parent checkpoints or training datasets—consistent.
-By comparing successive graph versions, our update mechanism applies only the incremental changes ($\Delta_{t}$) rather than rebuilding the full graph—keeping updates both scalable and auditable in practice.
-Figure 5. The temporal update for the datasets and different models on Hugging Face from June 25 to July 15, 2025.
-Finding #6: The $\Delta$-based update mechanism offers clear visibility into the LLM ecosystem’s rapid and continuous evolution, with daily growth primarily driven by fine-tuned, adapter, and quantized models.
-It enables efficient, audit-friendly tracking by applying only incremental changes, eliminating the need to rebuild the entire graph.
-These capabilities ensure that HuggingGraph remains both scalable and precise in capturing the dynamic nature of the LLM supply chain.
-5. Related Work
-This section aims to provide in-depth coverage of the current research landscape concerning the opportunities within the LLM supply chain, mainly emphasizing the different dimensions of the relationship between the models and datasets, focusing on the model and dataset hosting platforms (Sindhu et al., 2024; Hadi et al., 2024).
-LLM supply chain perspectives in AI.
-LLMs mark the revolutionary era in AI, with the edge passing through an exponential growth phase.
-A recent paper presents LLMs key components, including model infrastructure, lifecycle, and downstream applications (Wang et al., 2024).
-Another work showed that reusing these models has become so widespread, which encourages the sharing and adaptation of foundation models on a larger scale (Jiang et al., 2023).
-An open-source AI ecosystem, such as Hugging Face, hosts a broad range of LLMs and datasets and, therefore, plays an important role in democratizing AI technologies (Riva et al., 2024).
-The foundation models are key ingredients in such an ecosystem (Xu et al., 2024b).
-They can capture vast amounts of information from various datasets, allowing for effective task-specific variants like fine-tuning (Han et al., 2021).
-This trend underscores the significant impact of technological advancements on the democratization and innovation in AI (Cupać et al., 2024).
-Table 6. Daily $\Delta$ statistics for HuggingGraph updates (June 25–July 15, 2025), showing average additions, net changes, and peak values by node and edge type.
-Category
-New-
-per day
-Net Change-
-per day
-Peak-
-in month
-Datasets 39 $\pm$ 62 +785 849
-Base model 2 $\pm$ 5 +57 42
-Fine-tune 126 $\pm$ 168 +2526 2090
-Adapter 41 $\pm$ 64 +828 887
-Quantized 16 $\pm$ 19 +321 290
-Merge 1 $\pm$ 2 +34 13
-Dependency edges 23 $\pm$ 6 +189 289
-Relationship analysis between LLM models and datasets.
-A recent study investigates the practical adaptation of foundation models to specific domains and tasks.
-Multitask fine-tuning has demonstrated the potential to enhance performance on target tasks with scarce labels (Xu et al., 2024a).
-In plant phenotyping, adapting vision foundation models by techniques like adapter tuning and decoder tuning has shown results comparable to those of leading task-specific models (Chen et al., 2023).
-The Quadapter technique for language models tackles quantization difficulties by incorporating learnable parameters that scale activations channel-wise, mitigating overfitting during quantization-aware training (Park et al., 2022).
-The PathoTune framework in pathology imaging utilizes multi-modal prompt tuning to reconcile domain discrepancies among foundation models, tasks, and specific instances, exhibiting enhanced performance compared to single-modality methods and facilitating the direct adaptation of natural visual foundation models to pathological tasks (Lu et al., 2024a).
-These findings underscore the efficacy of various adaptation approaches in improving foundation model performance across multiple areas.
-From the above, we can see that models heavily depend on each other for fine-tuning, adaptation, or quantization.
-Building on these insights, two other works have indicated that serious security vulnerabilities may encumber the LLM supply chain to a few developers (Kathikar et al., 2023; Osborne et al., 2024).
-These factors may affect the diversity of innovation.
-The authors relate several challenges arising during the software engineering, security, and privacy of different relations and components involved in model creation and deployment (Huang et al., 2023; Wang et al., 2024).
-Moreover, the vulnerabilities can be transferable from one model to another model during the fine-tuning process (Hajipour et al., 2024).
-Therefore, serious security challenges are also posed because the vulnerabilities of the repositories threaten the whole ecosystem (Kathikar et al., 2023).
-Comparison with HuggingGraph.
-Our work is different from prior works in three aspects.
-First, the motivation for this work stems from the transferable property (Hajipour et al., 2024) of one model to others and during the training process with the datasets.
-Because of this, our analysis is mainly about the hierarchical relationship between the models, which model is used to fine-tune, adapt, or quantize other models, and which datasets are used to train the models.
-As it is challenging to get consistent hierarchical relational information, we made a supply chain graph from the metadata using APIs we got from the models and datasets.
-Second, as our analysis is one of the premier tasks in constructing a supply chain graph, for analyzing the relationship, we focus on the different graph properties like connectivity and WCC analysis.
-Third, we use forward and backward supply chain analysis to examine the built graph and learn more about the hierarchies of the models and datasets during fine-tuning, adaptation, quantization, and training.
-This is different from previous methods, which mostly focused on training the models.
-This would facilitate determining and identifying the relationship between the model and the dataset.
-More importantly, we conclude with multiple interesting findings, which we believe could provide important insights to the LLM community.
-6. Discussion
+In addition, models like static-similarity-mrl-mul-v1 and static-similarity-mrl-multilingual are both fine-tuned on 108 datasets, indicating their role in multilingual and multi-task similarity-based retrieval applications.
+The ModernBERT-base-embed and Llama-3.2-3B-Instruct families show strong dataset diversity as well, with 88 and 87 datasets, respectively, across fine-tuning and quantized variants (e.g., GGUF format).
+Interestingly, most of these models are fine-tuned, specifically, 8 out of the top 10, highlighting a trend where base models are transformed into task-specific variants through diverse training datasets.
+This pattern suggests that fine-tuning remains a dominant strategy for adapting base models to downstream tasks across heterogeneous data sources.
+Finding #5: Models and datasets exhibit strong bidirectional interdependence, with datasets like Mistral-7B-v0.1 spawning hundreds of models, while models such as DeBERTa-ST-AllLayers-v3.1 leverage diverse datasets to enhance adaptability, highlighting the critical role of dataset-model interactions in advancing AI.
+4.6. RQ #6: Dynamic Update Evaluation
+Figure 5. The number of changed models and datasets on Hugging Face from June 25 to July 15, 2025.
+This research question aims to understand the dynamic update of the LLM supply chain.
+Following the method discussed in Section 3.4, we perform a daily-based data collection by capturing how many nodes and edges are added and deleted each day.
+Figure 5 illustrates the sum of daily added and deleted of six key node categories (base, fine-tuned, adapters, quantized, merge variants, and datasets) from June 25 to July 15, 2025.
+We make three interesting observations.
+(i) The daily dynamic update is significant.
+That is, an average of 4,622 models are changing every day, including $\sim$3,843 model additions and $\sim$779 deletions.
+In addition, about 1,538 datasets are changing each day, containing $\sim$1,305 dataset additions and $\sim$233 deletions.
+An addition occurs when a new model or dataset is uploaded to the Hugging Face platform.
+This includes base models, task-specific variants, and new training datasets.
+A deletion refers to the removal of such nodes, often due to licensing issues, privacy concerns, or contributor decisions, such as replacing outdated models or withdrawing low-quality or sensitive datasets.
+(ii) Fine-tuned models dominate daily activity, averaging over 2,988 changes per day, followed by consistent contributions from adapters ($\sim$1,224/day) and datasets ($\sim$1,539/day).
+Noticeable spikes, such as on July 7 and July 9, align with major events like the Mistral-Fusion-v3 fine-tuning wave and dataset updates such as HFTime2025-News.
+(iii) Furthermore, adapter uploads peaked at 1,249 on June 28, while quantized variants reached 381 on July 9, driven by releases such as QWin-GGUF-7B.
+These patterns demonstrate HuggingGraph’s ability to capture evolving supply chain dynamics at a fine-grained level.
+Finding #6: The LLM supply chain exhibits continuous and high-volume daily changes, driven by frequent additions and deletions of models and datasets.
+This reflects a rapidly evolving and highly dynamic ecosystem shaped by active contributor behavior.
+4.7. RQ #7: Generability to Other AI Platforms
+To validate HuggingGraph’s generalizability beyond Hugging Face, we applied our pipeline to another AI platform, Kaggle (Bojer and Meldgaard, 2021).
+As of July 25, 2025, Kaggle hosts 470 base and task-specific models with 3,146 variants and approximately 502K datasets.
+Using Kaggle’s kernel and dataset APIs, we collected 2,640 models and 105,867 datasets.
+This significant gap is due to lack of the models’ and datasets’ metadata.
+Of the datasets retrieved ($\sim$105K), only 137 datasets were included in our graph, as most lacked standardized documentation or traceable links to models.
+Many are standalone, poorly described, or lack contextual information, a challenge also observed on Hugging Face.
+We follow the same way to construct a heterogeneous graph consisting of 2,777 nodes, which include 2,640 model nodes, comprising 59 base models, 2,410 fine-tuned, 171 quantization models, and 137 datasets.
+The graph contains a total of 3,990 edges, on which we observed seven types of edges, (i) base model $\rightarrow$ fine-tuned model (467 edges), (i) base model $\rightarrow$ quantization model (62 edges), (iii) fine-tuned model $\rightarrow$ fine-tuned model (1,696 edges), (iv) fine-tuned model $\rightarrow$ quantized model (107 edges), (v) quantized model $\rightarrow$ quantized model (1 edges), (vi) dataset $\rightarrow$ fine-tuned model (1,614 edges), and (vii) dataset $\rightarrow$ quantization model (43 edges).
+We observed that the average degree is 1.44, indicating that the graph is sparse.
+We made two interesting observations.
+(i) The degree distribution is heavy-tailed and skewed: out of 2,777 total nodes, 2,305 nodes ($\approx$83%) have a total degree of 1.
+Most nodes have low degrees, while a few highly connected hubs dominate the graph.
+For example, in-degrees range from 0 to 4 (with NaN peaking at 4), and tensorflow/mobilenet-v1 has the highest out-degree of 64, followed by google/nnlm with 56.
+(ii) Furthermore, the graph contains 448 weakly connected components (WCCs), reflecting high fragmentation.
+However, the largest WCC includes 65 nodes, suggesting the presence of a moderately sized core subgraph.
+Finding #7: The resulting graph exhibits structural properties consistent with our Hugging Face analysis, including a heavy-tailed degree distribution, sparse connectivity, and strong modular fragmentation, demonstrating the robustness and generalizability of our pipeline across platforms despite metadata limitations.
+5. Use Case
 HuggingGraph presents a technique to analyze the supply chain of the LLM ecosystem.
@@
 HuggingGraph can be used to address this challenge by constructing the supply chain of models and datasets, uncovering both direct and derived dependencies, even when they are not formally documented.
-For example, it can trace how the model Meta-llama indirectly relies on a dataset like Wikimedia via Awesome-Chatgpt-prompt (Figure 2).
-This transparency supports developers, auditors, and policymakers in validating provenance, detecting risks, and enabling responsible, trustworthy AI deployment.
+For example, it can trace how the model Meta-llama indirectly relies on a dataset like Wikimedia via Chatgpt-prompt (Figure 2).
+This transparency supports developers, auditors, and policymakers in validating provenance, detecting risks, and enabling trustworthy AI.
 Use case #2: Identifying critical nodes and structural vulnerabilities.
@@
 HuggingGraph can be used to address this by modeling the supply chain as a graph and analyzing node connectivity to surface highly reused models and datasets with significant inbound or outbound links.
-This visibility enables maintainers to safeguard vital assets and helps developers mitigate the risk of overreliance on fragile or under-maintained components, fostering more robust and secure model development pipelines.
+This visibility enables maintainers to safeguard vital assets and helps developers mitigate the risk of overreliance on fragile or under-maintained components.
+6. Related Work
+This section aims to provide in-depth coverage of the current research landscape concerning the opportunities within the LLM supply chain, mainly emphasizing the different dimensions of the relationship between the models and datasets, focusing on the model and dataset hosting platforms (Sindhu et al., 2024; Hadi et al., 2024).
+LLM supply chain perspectives in AI.
+LLMs mark the revolutionary era in AI, with the edge passing through an exponential growth phase.
+A recent paper presents LLMs key components, including model infrastructure, lifecycle, and downstream applications (Wang et al., 2024).
+Another work showed that reusing these models has become widespread, which encourages the sharing and adaptation of base models on a larger scale (Jiang et al., 2023).
+An open-source AI ecosystem, such as Hugging Face, hosts a broad range of LLMs and datasets and, therefore, plays an important role in democratizing AI technologies (Riva et al., 2024).
+The base models are key ingredients in such an ecosystem (Xu et al., 2024b).
+They capture vast amounts of information from various datasets, allowing for effective task-specific variants like fine-tuning (Han et al., 2021).
+This trend underscores the significant impact of technological advancements on AI democratization and innovation (Cupać et al., 2024).
+Relationship analysis between LLM models and datasets.
+A recent study investigates the practical adaptation of base models to specific areas and tasks.
+Multitask fine-tuning has demonstrated the potential to enhance performance on target tasks with scarce labels (Xu et al., 2024a).
+In plant phenotyping, adapting vision-based models by techniques like adapter tuning and decoder tuning has shown results comparable to those of leading task-specific models (Chen et al., 2023).
+The Quadapter technique for language models tackles quantization difficulties by incorporating learnable parameters that scale activations channel-wise, mitigating overfitting during quantization-aware training (Park et al., 2022).
+These findings underscore the efficacy of various adaptation approaches in improving base model performance across areas.
+From the above, we can see that models heavily depend on each other for fine-tuning, adaptation, quantization or merger.
+Building on these insights, two other works have indicated that serious security vulnerabilities may encumber the LLM supply chain to a few developers (Kathikar et al., 2023; Osborne et al., 2024).
+These factors may affect the diversity of innovation.
+The authors relate several challenges arising during the software engineering, security, and privacy of different relations and components involved in model creation and deployment (Huang et al., 2023; Wang et al., 2024).
+Moreover, the vulnerabilities can be transferable from one model to another model during the fine-tuning process (Hajipour et al., 2024).
 7. Conclusion
@@
 First, we design a methodology to systematically collect the supply chain information of LLMs.
-With the collected information, we design a new graph to model their relationships, which is a large directed heterogeneous graph, having 397,376 nodes and 453,469 edges.
-From there, we perform different types of analysis and make multiple interesting findings.
+With that, we construct a large directed heterogeneous graph, having 402,654 nodes and 453,469 edges.
+Then, we perform different types of analysis and make multiple interesting findings.
 Acknowledgment

### v3

@@
 Abstract.
-Large language models (LLMs) leverage deep learning architectures to process and predict sequences of words based on context, enabling them to perform a wide range of natural language processing tasks, such as translation, summarization, question answering, and content generation.
-However, the increasing size and complexity of developing, training, and deploying cutting-edge LLMs demand extensive computational resources and large-scale datasets.
-This creates a significant barrier for researchers and practitioners.
-Because of that, platforms that host models and datasets have gained widespread popularity.
-For example, on one of the most popular platforms, i.e., Hugging Face, there are more than 1.8 million models and more than 450K datasets by the end of June 2025, and the trend does not show any slowdown.
+Large language models (LLMs) leverage deep learning architectures to process and predict sequences of words, enabling them to perform a wide range of natural language processing tasks, such as translation, summarization, question answering, and content generation.
 As existing LLMs are often built from base models or other pre-trained models and use external datasets, they can inevitably inherit vulnerabilities, biases, or malicious components that exist in previous models or datasets.
-Therefore, it is critical to understand these components’ origin and development process to detect potential risks better, improve model fairness, and ensure compliance with regulatory frameworks.
+Therefore, it is critical to understand these components’ origin and development process to detect potential risks, improve model fairness, and ensure compliance with regulatory frameworks.
 Motivated by that, this project aims to study such relationships between models and datasets, which are the central parts of the LLM supply chain.
-First, we design a methodology to collect LLMs’ supply chain information systematically.
-With the collected information, we design a new graph to model the relationships between models and datasets, which is a large directed heterogeneous graph, having 402,654 nodes and 462,524 edges.
-Then, on top of this graph, we perform different types of analysis and make multiple interesting findings.
+First, we design a methodology to systematically collect LLMs’ supply chain information.
+Then, we design a new graph to model the relationships between models and datasets, which is a directed heterogeneous graph, having 402,654 nodes and 462,524 edges.
+Lastly, we perform different types of analysis and make multiple interesting findings.
+Large language models (LLMs); AI supply chain; Graph analysis
+journalyear: 2025copyright: ccconference: Proceedings of the 34th ACM International Conference on Information and Knowledge Management; November 10–14, 2025; Seoul, Republic of Koreabooktitle: Proceedings of the 34th ACM International Conference on Information and Knowledge Management (CIKM ’25), November 10–14, 2025, Seoul, Republic of Koreadoi: 10.1145/3746252.3761510isbn: 979-8-4007-2040-6/2025/11ccs: Computing methodologies Artificial intelligenceccs: Theory of computation Graph algorithms analysisccs: Applied computing Supply chain management
 1. Introduction
 Large language models (LLMs) are AI models designed to understand and generate human language by learning patterns and relationships within extensive datasets (Mirchandani et al., 2023; Shen et al., 2023), such as GPT (Generative Pre-trained Transformer) (Yenduri et al., 2024), BERT (Bidirectional Encoder Representations from Transformers) (Kenton and Toutanova, 2019), and T5 (Text-To-Text Transfer Transformer) (Bahani et al., 2023).
-These models leverage deep learning architectures to process and predict sequences of words based on context, enabling them to perform a wide range of natural language processing tasks (Bonner et al., 2023), such as, translation (Lu et al., 2024), summarization (Laban et al., 2023), question-answering (Allemang and Sequeda, 2024), and content generation (Agossah et al., 2023).
-LLMs usually have billions (or even trillions) of parameters (Merrick et al., 2024), enabling them to generate high-quality text.
+These models leverage deep learning architectures to process and predict sequences of words based on context, enabling them to perform a wide range of tasks (Bonner et al., 2023), such as, translation (Lu et al., 2024), summarization (Laban et al., 2023), question-answering (Allemang and Sequeda, 2024), and content generation (Agossah et al., 2023).
+LLMs usually have billions ore more parameters (Merrick et al., 2024), enabling them to generate high-quality text.
+Figure 1. The number of AI models and datasets (in million scale) on Hugging Face from July 2024 to June 2025.
 However, the increasing size and complexity of developing, training, and deploying cutting-edge LLMs demand extensive computational resources (Xu et al., 2024b) and large-scale datasets (Villalobos et al., 2024).
 This creates a significant barrier for researchers and practitioners, limiting their access to state-of-the-art models (Mizrahi et al., 2024).
-As the demand for democratizing access to such LLM models continues to rise, platforms that host models and datasets have gained widespread popularity.
-For example, Figure 1 shows the number of models and datasets (in a million scales) on Hugging Face, the largest public AI model hosting platform (Face, [n.
-d.]), starting from July 2024 to June 2025.
+As the demand for democratizing access to such LLM models continues to rise, platforms that host models and datasets have gained widespread popularity, e.g., Hugging Face (Face, [n.
+d.]), ONNX Model Zoo (ONNX Community, 2025), and PyTorch Hub (PyTorch Core Team, 2024).
+Figure 1 shows the number of AI models and datasets (in million scale) on Hugging Face, one of the largest public AI model hosting platforms (Face, [n.
+d.]), from July 2024 to June 2025.
 By the end of June 2025, it has reached over 1.8M models and 450K datasets.
 In addition, the trend does not show any slowdown.
-Such LLM hosting platforms provide user-friendly interfaces, APIs, and cloud-based infrastructures that enable researchers and developers to easily share, fine-tune, and deploy models without requiring extensive computational resources.
-Moreover, they foster open collaboration, allowing the broader community to contribute to model improvements, benchmark performances, and enhance transparency in AI research.
-On such platforms, different types of models can be classified into two categories based on their tasks, i.e., base models and task-specific models.
+Such platforms provide user-friendly interfaces, APIs, and cloud-based infrastructures that enable researchers and developers to easily share, fine-tune, and deploy models without requiring extensive computational resources.
+Based on the tasks, these models can be classified into two broad categories, base models and task-specific models.
 (i) Base models are large, pre-trained models that can be fine-tuned for specific downstream tasks (Tan et al., 2024).
@@
 First, fine-tuned models adapt base models for specific tasks by training on additional task-specific datasets (Zou et al., 2023).
-Second, adapters models add lightweight and modular layers to the pre-trained models for specific tasks (Hu et al., 2023).
+Second, adapter models add lightweight and modular layers to the pre-trained models for specific tasks (Hu et al., 2023).
 Third, quantization models trade off precision in numerical computations for accelerating inference and reducing memory consumption (e.g., using less precise model parameters) (Velingker et al., [n.
 d.]).
-Fourth, merged models integrate multiple models into a single unified model by combining weights or configurations enabling support for multiple tasks or domains without separate deployments (Akiba et al., 2025).
-Besides models, such platforms also host many datasets used for training and fine-tuning the previously discussed models (Röttger et al., 2024).
+Fourth, merged models integrate multiple models into a single unified model by combining weights or configurations (Akiba et al., 2025).
+Besides, such platforms also host many datasets used for training or adapting (e.g., fine-tuning) models (Röttger et al., 2024).
 1.1. Motivation
-As existing LLMs are often built from base models or other pre-trained models and using external datasets, they can inevitably inherit vulnerabilities, biases, or malicious components from previous models or datasets.
+As existing LLMs are often built from base models or other pre-trained models and use external datasets, they can inevitably inherit vulnerabilities, biases, or malicious components from previous models or datasets.
 Thus, understanding these components’ origin and provenance can help better detect potential risks, improve model fairness, and ensure compliance with regulatory frameworks.
-Figure 1. The number of AI models and datasets (in a million scale) on Hugging Face from July 2024 to June 2025.
 Motivated by that, this paper aims to study such relationships between models and datasets.
@@
 With that, one can easily keep track of them.
-For example, a use case is when a security vulnerability is found in one LLM, and we can quickly locate the potential models that might have the same issues.
+For example, a use case is when a security vulnerability is found in one LLM, and we can quickly locate the potential models that might have the same vulnerabilities.
 Dataset origin.
-This supply chain can also help to understand the datasets’ origins used for training different models (Shen et al., 2023).
+This supply chain can also help to understand dataset origin used for training different models (Shen et al., 2023).
 Dataset origin refers to the source from which the data is collected.
@@
 In this paper, we mainly study the most popular AI platform, i.e., Hugging Face, but the same strategy applies to other platforms.
-In particular, we use the APIs from the AI platform to collect the metadata about the hosted model and dataset.
+In particular, we use the APIs from the AI platform to collect the metadata about the hosted models and datasets.
 To this end, we collected a large dataset as of June 30, 2025.
-Second, with the collected metadata, we created a new graph, named LLM supply chain graph, to model the relationships between models and datasets.
-It is a directed heterogeneous graph where a node denotes different types of datasets and models (including base, fine-tune, adapter, quantization, and merge).
-An edge denotes the dependency relationship between them, including model-model, dataset-dataset, and model-dataset relationships.
+Second, with the collected metadata, we construct a new graph, named LLM supply chain graph [^1] The constructed graph: https://github.com/SC-Lab-Go/HuggingGraph [^2] A demonstration website: https://ai-supply-chain.github.io/, to model the relationships between models and datasets.
+It is a directed heterogeneous graph, where a node denotes different types of models and datasets, and an edge denotes the dependency between them.
 Together, this graph is able to accurately capture the LLM supply chain information.
-To this end, we constructed a large graph with 402,654 nodes and 462,524 edges.
-An anonymized version of the complete graph is publicly available at GitHub [^1] LLM supply chain graph: https://github.com/huggingface00/HuggingGraph.
-Third, with this graph, we perform different types of analysis, including forward and backward analysis.
-We study seven research questions, including (i) the properties of the LLM supply chain graph, (ii) structural analysis, (iii) supply chain relationships between AI models, (iv) supply chain relationships between datasets, (v) supply chain relationships between models and datasets, (vi) dynamic update evaluation, and (vii) extension to other AI platforms.
+To this end, we get a graph with 402,654 nodes and 462,524 edges.
+Lastly, we perform different types of analysis, including forward and backward analysis.
+We study six research questions, including (i) the properties of the LLM supply chain graph, (ii) structural analysis, (iii) supply chain relationships between AI models, (iv) supply chain relationships between models and datasets, (v) dynamic update evaluation, and (vi) generalizability to other AI platforms.
+We hope this study can not only provide insights on LLM supply chain, but also raise awareness and future interests in this direction.
 2. Preliminary
-The LLM supply chain encompasses the interconnected processes required for developing, deploying, and maintaining models (Wang et al., 2024).
+LLM supply chain encompasses the interconnected processes required for developing, deploying, and maintaining models (Wang et al., 2024).
 This includes sourcing and preparing data to ensure high-quality and diverse datasets (Wang et al., 2024).
-It also involves creating and training models (Li et al., 2023).
-Finally, it covers making trained models available through APIs (Singla et al., 2023).
-In addition, LLMs can undergo adaptation, quantization, and fine-tuning, a process where they are tuned with domain-specific datasets to maximize performance on specific tasks (VM et al., 2024), thus improving their accuracy and applicability.
+It also involves creating and training models (Li et al., 2023), and making trained models available through APIs (Singla et al., 2023).
+In addition, LLMs can undergo fine-tuning, adaptation, quantization, and merging processes in which they are tuned with domain-specific datasets to maximize performance on specific tasks (VM et al., 2024), thus improving their accuracy and applicability.
 This study mainly focuses on the relationships between models and datasets, which are the central parts of the whole LLM supply chain ecosystem.
-We hope this study can not only provide insights on the LLM supply chain but also raise awareness and future research interests in this direction.
 3. Methodology
@@
 To analyze the LLM supply chain ecosystem, we need a large dataset with such information.
-Fortunately, platforms like Hugging Face provide some APIs that allow us to access the model and dataset and collect their metadata information, which can be used to construct the supply chain ecosystem.
+Fortunately, platforms like Hugging Face provide some APIs that allow us to access the model and dataset and collect their metadata, which can be used to construct the LLM supply chain graph.
 Table 1 summarizes the four types of APIs we used.
@@
 API Name Description
-Model hub Access the model hub to list, search, and download models and metadata.
-Dataset Access the datasets for discovery, metadata retrieval, and downloading.
-Metrics Access metrics for model evaluation, e.g., metric discovery, metadata retrieval, and calculation.
+Model hub Access model hub to list, search, and download models and metadata.
+Datasets Access the datasets for discovery, metadata retrieval, and downloading.
+Metrics Access metrics for model evaluation, e.g., metric discovery.
 Search Search name, tag, or other metadata for model and dataset.
 Handling missing information.
-Accurate construction of the LLM supply chain graph critically depends on the quality of metadata from the LLM platforms (e.g., Hugging Face), which might suffer from missing or incomplete data.
-To address this limitation, we apply the following two techniques, i.e., cross-reference links, and textual pattern extraction.
+Accurate construction of the LLM supply chain graph depends on the quality of metadata from the LLM platforms (e.g., Hugging Face), which might suffer from missing or incomplete data.
+To address it, we design two techniques, i.e., cross-reference links, and textual pattern extraction.
 (i) Cross-reference links.
@@
 3.2. LLM Supply Chain Graph
-With the collected metadata, we construct a directed heterogeneous graph to model the LLM supply chain.
-In this graph, a node denotes different types of datasets and models, including base, fine-tune, adapter, quantization, and merge models.
+The collected metadata can be accurately modeled by the graph data structure.
+It is a directed heterogeneous graph, where a node denotes different types of datasets and models, including base, fine-tune, adapter, quantization, and merge models.
 An edge denotes the dependency relationship between them, including model-model, dataset-dataset, and model-dataset relationships.
@@
 As shown in Figure 2, model Llama-3.3-70B is fine-tuned from the base model Meta-llama.
-Then, it is used by the models Doctor-Shotgun and Llama-3.3-70B-4bit-Vision to generate an adapter and quantization model, respectively.
+Then, it is used by the models Doctor-Shotgun and Llama-3.3-70B-4bit to generate an adapter and quantization model, respectively.
 (ii) Dataset–dataset relationship.
@@
 (iii) Model-dataset relationship.
-To capture the model and dataset dependency, we use the metadata from both models and datasets.
+To capture this, we use the metadata from both models and datasets.
 The metadata of a model might specify the datasets used for training or adapting.
 However, not all the models disclose such information.
-To capture more information, we find the metadata of a dataset contains a data field of “trained_fine_tune_models” on this dataset.
-Thanks to that, we are able to capture the accurate model and dataset relationships.
+Fortunately, we find the metadata of a dataset contains a field of “trained_fine_tune_models” on this dataset.
+Thanks to that, we can capture the accurate model and dataset relationship.
 In Figure 2, the datasets The Pile and Chatgpt-prompt have directed edges to model Meta-Llama, meaning that both datasets are used to train the model.
@@
 This supply chain graph can help to understand the transformational processes of the models and datasets.
-In particular, we can understand how base models evolve into their variants, including fine-tuned, adaptive, quantized, or merged models, and vice versa.
+In particular, we can understand how base models evolve into their variants, including fine-tuned, adapted, quantized, or merged models, and vice versa.
 Similar observations can be made for datasets.
@@
 Forward analysis is the method of traversing the supply chain graph following the dependency edges of a chosen node in a forward-going way.
-This node (known as the root/source node) can be a dataset, a base model, a fine-tuned model, an adapter, or a quantized model.
-In particular, given a source node, we apply the graph traversal algorithm (e.g., breadth-first search (BFS) (Tarjan, 1972)) to traverse all the nodes (including both models and datasets) in a level-by-level pattern.
+This node (known as the root/source node) can be a dataset, a base model, a fine-tuned model, an adapter, a quantized, or a merged model.
+In particular, given a source node, we apply the graph traversal algorithm (e.g., breadth-first search (BFS) (Knuth, 1974)) to traverse all the nodes (including both models and datasets) in a level-by-level pattern.
 To that end, this forward analysis will identify all the nodes that are reachable from the source node.
 Model analysis example.
-In Figure 2, to analyze the forward supply chain of Meta-Llama, we trace its development from training datasets to subsequent adaptations.
-This model is fine-tuned into Llama-3.3-70B, enhancing its capabilities for specific tasks.
-Summarizing the forward supply chain of Meta-Llama, we identify the following four distinct forward supply chains: (i) Base model (Meta-llama) $\rightarrow$ fine-tuned model (Llama-3.3-70B) $\rightarrow$ another fine-tuned model (Llama3.3-70B-Vision).
+In Figure 2, we analyze the forward supply chain of Meta-Llama.
+In particular, we identify four distinct forward paths: (i) Base model (Meta-llama) $\rightarrow$ fine-tuned model (Llama-3.3-70B) $\rightarrow$ another fine-tuned model (Llama3.3-70B-Vision).
 (ii) Base model (Meta-llama) $\rightarrow$ adapted model (Doctor-Shotgun).
@@
 (iv) Base model (Meta-llama)) $\rightarrow$ merged model (MistLlama).
-These pathways illustrate the evolution trajectory of the base model Meta-Llama showcasing its progressive specialization and adaptation for various tasks.
+These paths show the evolution trajectory of the base model Meta-Llama, showcasing its progressive specialization and adaptation for various tasks.
 Dataset analysis example.
@@
 Together, these datasets form a unified corpus that serves as training data for models like Meta-LlaMA.
-Backward analysis, on the other hand, is the method of traversing the supply chain graph following the edges in a backward way.
-We accomplish this by traversing the directed graph also with BFS (Knuth, 1974) in reverse, starting from the selected node and following the incoming edges.
+Backward analysis is the method of traversing the supply chain graph following the edges in a backward way.
+We accomplish this by traversing the graph also with BFS (Knuth, 1974), starting from the selected node and following the incoming edges.
 To that end, this backward analysis will identify all the nodes that can reach the source node.
@@
 This model is fine-tuned from Unsloth, which in turn originates from its base model, Meta-Llama, and the datasets used to train the base model are The Pile and Chatgpt-prompts.
-Through this analysis, we establish the backward path, starting from the target model, i.e., RBot70Bv4, and tracing to its base model, Meta-Llama, revealing dependencies and transformations involved in its development.
-Similarly, we can analyze datasets.
+Through this analysis, we establish the backward path, starting from target model, RBot70Bv4, and tracing to its base model, Meta-Llama, revealing dependencies and transformations in its development.
+Similar methods can be used for dataset analysis.
 3.4. Accommodating Dynamic Update
+Figure 3. (a) Indegree and (b) outdegree distribution. X-axis shows degrees, and Y-axis shows node count (log scale).
 The hosted models and datasets on AI platforms are growing fast as new models are being developed every day.
-For example, between June 25 and July 15, 2025, we observed 80,703 new models (approximately 3,843 per day), 27,405 new datasets (approximately 1,305 per day) on Hugging Face only, which is just one of the many AI platforms.
+For example, between June 25 and July 15, 2025, we observed 80,703 new models (approximately 3,843 per day), 27,405 new datasets (approximately 1,305 per day) on Hugging Face, which is just one of the many AI platforms.
 Therefore, we need to accommodate the dynamic update to accurately manage and analyze the AI supply chain.
-Particularly, HuggingGraph accommodates the dynamic update in three steps.
+Particularly, HuggingGraph accommodates the dynamic updates in three steps.
 (i) Scoping updated models or datasets.
 At a time $t$, we keep a copy of the hosted models and datasets with their IDs.
-When we evolved to $t+1$, we get another copy of the hosted models and datasets with their IDs.
+When it evolved to $t+1$, we get another copy of the hosted models and datasets with their IDs.
 The difference between them shows the updated models and datasets, including newly added or deleted.
@@
 Given the newly updated dependency $\Delta_{t+1}$, and let $G_{t}$, $G_{t+1}$ denote the graph at time $t$, $t+1$, respectively, then $G_{t+1}=G_{t}\cup\Delta_{t+1}$.
-4. Experiments and Finding
-To deeply understand the relationships between models and datasets, we study seven critical research questions (RQs), which can offer valuable insights into the supply chain of the LLM ecosystem and potentially pave the way for future LLM development.
+4. Experiment and Finding
+To deeply understand the relationships between models and datasets, we study six critical research questions (RQs) as below.
   • RQ #1: What are the properties of LLM supply chain graph?
@@
   • RQ #3: What are the supply chain between LLM models?
-  • RQ #4: What are the supply chain between datasets?
-  • RQ #5: What are the relationships between models and datasets?
-  • RQ #6: What insights can be gained from the dynamic updates?
-  • RQ #7: How can HuggingGraph be applied to other platforms?
+  • RQ #4: What are the relationships between models and datasets?
+  • RQ #5: What insights can be gained from the dynamic updates?
+  • RQ #6: How can HuggingGraph be applied to other platforms?
 4.1. RQ #1: Supply Chain Graph Properties
-Figure 3. (a) Indegree distribution and (b) Outdegree distribution. X-axis represents degrees, and the Y-axis represents the number of vertices in a logarithmic scale.
 This research question aims to understand the critical properties of LLM supply chain graph, i.e., graph basics and degree distribution.
@@
 Furthermore, we identified substantial metadata missing.
-As of June 30, 2025, among 1.8 million models, only 50,156 (2.76%) include a model tree, while $\sim$ 550K models lack any metadata beyond their name, with nearly 400K entries empty or invalid.
-Similarly, of the 450K datasets, only 68,634 (15.26%) provide datacards, leaving $\sim$ 380K datasets without any usable metadata.
+As of June 30, 2025, among 1.8 million models, only 50,156 (2.79%) provides a model tree, while $\sim$550K models lack any metadata beyond their names, and nearly another 400K models are empty.
+Similarly, of the 450K datasets, only 68,634 (15.26%) provide dataset cards, leaving $\sim$380K datasets without any metadata.
 This highlights a broader issue in the AI community, where a significant number of models and datasets lack consistent and structured documentation on the supply chain.
 This reflects the need of more transparent disclosure.
-Degree distribution of a graph describes how node degrees (the number of edges connected to a node) are distributed across the graph.
+Degree distribution shows how node degrees (the number of edges connected to a node) are distributed across the graph.
 Figure 3 illustrates the indegree and outdegree distribution of the graph.
 We show not only the total distribution but also the distribution of six types of nodes, including base models, fine-tuned models, adapter models, quantization models, merged models, and datasets.
-We made two interesting observations.
-(i) This degree distribution in our supply chain graph shows a heavy-tailed behavior.
+We observe that the degree distribution in our supply chain graph is heavy-tailed.
 In particular, the indegree distribution shows a large spread across different categories.
 The outdegree distribution follows a similar pattern but may differ in specific cases (e.g., adapters seem to have a more restricted degree distribution).
-The heavy-tailed behavior suggests that most nodes have low degrees, while a few central nodes (hubs) dominate the graph.
-In particular, the dataset macrocosm-os/images has the highest indegree value of 550, and Mistral-7B-v0.1 from “mistral AI” has the highest outdegree value of 1,093.
+This heavy-tailed behavior suggests that most nodes have low degrees, while a few central nodes (hubs) dominate the graph.
+In particular, the dataset macrocosm-os/images has the highest indegree 550, and dataset Mistral-v0.1 from “mistral AI” has the highest outdegree 1,093.
 Specifically, the base models act as high-degree hub nodes as they are heavily used by other task-specific models.
@@
 4.2. RQ #2: Supply Chain Structural Analysis
-Figure 4. Cumulative distribution function (CDF) of WCC size.
+Figure 4. Cumulative distribution function (CDF) of WCC.
 This research question aims to understand the topology and evolution of the LLM supply chain. To achieve that, we analyze the structural properties with connectivity and community analysis.
-Connectivity analysis analyzes the structural connectivity information.
-We performed weakly connected components (WCC), which identifies maximal sets of nodes that remain connected when edge directions are ignored.
+Connectivity analysis.
+We computed weakly connected components (WCCs), which identify a maximal subset of nodes that remain connected when edge directions are ignored (Ji and Huang, 2020; Ji et al., 2020).
 The total number of WCCs in our supply chain graph is 44,908.
 Figure 4 shows the cumulative distribution function (CDF) of the WCC distribution.
-We made two interesting observations.
-(i) The largest WCC covers 247,244 nodes, accounting for 61.8% of all the nodes.
+We observe (i) the largest WCC covers 247,244 nodes, accounting for 61.4% of all the nodes.
 It reflects the dense interconnections that pervade the ecosystem.
-This vital element is essential for effective information sharing, resource allocation, and structural support and is the base of the ecosystem.
+This vital element is essential for effective information sharing, resource allocation, and structural support, and is the base of the ecosystem.
 In the largest WCC, major models are included, such as Gemma-2B, DistilBERT, and GPT-2.
-(ii) In contrast, the remaining WCCs collectively hold 38.6% of the nodes (i,e,.155,410 nodes), with most having 1, 2 or 3 nodes.
+(ii) In contrast, the remaining WCCs collectively hold 38.6% of the nodes, with most having 1, 2, or 3 nodes.
 This indicates a fragmented outer edge characterized by specialized models, rare datasets, or active experimental projects.
-The prevalence of these small, isolated pieces suggests niche attempts that lack integration with the overall system.
+The prevalence of these small, isolated pieces suggest niche attempts that lack integration with the overall system.
 For example, zhongqy/RMCBench (benchmarking dataset) and yigit69/bert-base-uncased-finetuned-rte-run_3 (recognizing textual entailment task) remain disconnected due to limited reuse or insufficient metadata.
-We also computed strongly connected component (SCC), a maximal subgraph in which every node is reachable from every other via directed edges, identifying 398,198 SCCs.
-Remarkably, only 591 of these are non‑trivial (size ¿ 1), collectively encompassing 2,169 nodes (0.54% of the graph).
+We also computed strongly connected component (SCC) (Ji et al., 2018), a maximal subgraph in which every node is reachable from every other via directed edges, identifying 398,198 SCCs.
+Remarkably, only 591 of these are non‑trivial (size ¿ 1), collectively encompassing 5,047 nodes (1.25% of the graph).
 The largest SCC comprises 478 dataset nodes, among them tree‑of‑knowledge and OpenHermes‑2.5, forming a tightly‑knit cluster.
 By contrast, the remaining 99.46% of nodes each reside in trivial (size‑1) SCCs.
-Community detection is the process of identifying groups or clusters of nodes within a graph that are more densely connected internally than with the rest of the graph.
-In the context of the LLM supply chain, it reveals semantically or functionally aligned subgraphs that reflect the patterns of reuse and task-specific models or datasets.
-We use the Louvain method (De Meo et al., 2011), which is a greedy optimization method for maximizing modularity.
-Modularity is a measure of how well a graph is divided into communities, where a high modularity score indicates that most edges fall within communities rather than between them.
-Table 2 summarizes the top-10 communities detected using the Louvain method.
-Our analysis reveals an exceptionally modular structure, with each of the top communities achieving a high modularity score of 0.96, indicative of strong intra-community connectivity.
+Community detection identifies node groups with dense internal connections.
+In LLM supply chains, it reveals aligned subgraphs reflecting reuse patterns and task-specific assets.
+We apply the Louvain method (De Meo et al., 2011), a greedy algorithm that maximizes modularity to find densely connected communities.
+Higher modularity indicates stronger intra-community connectivity.
+Table 2 summarizes the top-10 communities.
+(i) We find that each of the top-10 communities achieve a high modularity score of 0.96, indicating strong intra-community connectivity.
 Collectively, these communities span a wide range of functional domains, underscoring the presence of well-defined, task-aligned clusters within the ecosystem.
-We made two interesting observations.
-(i) The largest community consists of 9,390 nodes and attains a modularity score of 0.96, indicating an extremely cohesive internal structure.
+(ii) The largest community consists of 9,390 nodes and attains a modularity score of 0.96, indicating an extremely cohesive internal structure.
 It revolves around base models like OLMoE and CausalLM, and general-purpose datasets like prompt-perfect.
 This suggests a densely connected cluster facilitating widespread reuse and fine-tuning.
-(ii) Several other communities reflect clear task-based segmentation.
-For instance, Community 3 (7,388 nodes, modularity 0.96) focuses on benchmarking, with models like Wanxiang and datasets like marco-o1.
-Similarly, ID 5 (4,554 nodes, modularity 0.96) centers on instruction-tuning, connecting models like MedLlama-3-8B with curated datasets like dpo-mix.
+(iii) Several other communities reflect clear task-based segmentation.
+For instance, community 2 (7,388 nodes) focuses on solving mathematical problems, with models like qwen2.5_math and datasets like Marco-o1.
+Similarly, community 6 (4,554 nodes) centers on instruction-tuning, with model MedLlama-3-8B and dataset dpo-mix.
 Table 2. Top-10 Louvain communities sorted by size.
@@
 10 3,828 Mistral vicuna_format 0.96
-Finding #2: The LLM supply chain graph features a dominant core (61.8% of nodes), while high modularity (0.96) reveals task-aligned, semantically coherent communities amid a fragmented periphery.
-4.3. RQ #3: Supply Chain Analysis of LLM Models
+Finding #2: The LLM supply chain graph features a dominant core (61.4% of nodes), while high modularity (0.96) reveals task-aligned, semantically coherent communities amid a fragmented periphery.
+4.3. RQ #3: Supply Chain Analysis of Models
 This research question aims to provide a holistic view of the dependencies between the models within the LLM supply chain, particularly from both base and task-specific models.
@@
 The more dependencies, the larger the impact it has.
-We start with a base model and perform forward analysis by computing a breadth-first search (BFS) following the outgoing edges to get that from our LLM supply chain graph.
+We start with a base model and perform forward analysis by computing BFS following the outgoing edges.
 This leads to a forward subgraph, which denotes all the models that depend on the base model, including fine-tuned, adapted, quantized, or merged models.
 Table 3 shows the top-10 base models sorted by the forward subgraph size, which is the number of impacted task-specific models.
-We make two interesting observations.
-(i) A base model can significantly impact the LLM supply chain ecosystem.
-For example, Llama-3.1-8B is a base model from Meta used for efficient text generation, code assistance, and research (He et al., 2024).
+We find (i) a base model can significantly impact the LLM supply chain ecosystem.
+For example, Llama-3.1-8B is a base model from Meta used for efficient text generation, code assistance, and research (He et al., 2024b).
 Due to its relatively small size, which allows for deployment in resource-constrained environments, making advanced AI accessible to broader stakeholders (Grattafiori et al., 2024).
 It has generated up to 7,544 models, including 1,710 fine-tuned versions, 1,542 adapters, 3,473 quantizations, and 1,693 merged models tailored to specific tasks.
-(ii) In terms of fine-tuning, the base model Mistral-7B-v0.1 has been fine-tuned the most times, totaling 2,105.
-It is a faster and lighter version of the Mistral model from Mistral AI.
-It is obtained by training on a large corpus with grouped-query and sliding window attention by Mistral AI, delivering efficient text generation, NLP, and code assistance on consumer hardware for low-latency applications like chatbots and text classification (Thakkar and Manimaran, 2023).
+(ii) For fine-tuning, the base model Mistral-7B-v0.1 has been fine-tuned the most, totaling 2,105.
+It is a faster, lighter Mistral model trained by Mistral AI with grouped-query and sliding-window attention, enabling efficient text generation, NLP, and code assistance on consumer hardware for low-latency tasks (Thakkar and Manimaran, 2023).
 Table 3. Top-10 base models sorted by forward subgraph size.
@@
 Merge Level
-Llama-3.1-8B 7544 1710 1542 3473 1693 25
-Mistral-7B-v0.1 6744 2105 2187 1435 1254 27
-Qwen2.5-7B 6733 1972 1764 2516 1132 11
-Meta-Llama-3-8B 5633 967 1511 2220 1967 21
-Llama-3.1-70B 4063 698 281 2075 2519 11
-Qwen2.5-32B 3909 1086 158 2311 1049 12
-Qwen2.5-1.5B 3645 1300 1290 949 248 8
-Qwen2.5-0.5B 3521 1669 1006 810 46 11
-Qwen2.5-14B 3362 726 411 1880 1166 15
-Meta-Llama-3-8B-Instruct 3118 640 405 1394 1305 34
-Task-specific model analysis.
-Given a task-specific model, we want to understand how it evolves, meaning what other models it has relied on.
-To achieve that, for each task-specific model, we perform a backward analysis by running a BFS following the incoming edges.
-To that end, the derived subgraph shows the models it relies on.
-Table 4 shows the top 10 task-specific models sorted by the backward subgraph size, which is the number of models they rely on.
-We make two interesting observations.
-(i) A fine-tuned model, command-r-1-layer, illustrates the depth and complexity of transformations in the LLM supply chain.
-This model operates in bfloat16 (BF16) precision for efficient text generation and natural language understanding (Pignatelli et al., 2024), originates from the base model cc4ai, and has undergone extensive lineage evolution before reaching its final form.
-Specifically, it depends on 40 upstream artifacts, including 39 other fine-tuned models, and spans 39 transformation levels in its backward lineage chain, as detailed in Table 4.
-(ii) We observed that adapters are mainly used for lightweight fine-tuning and merges for model integration, but task-specific models like command-r-1-layer, as optimized standalone derivatives, do not evolve from adapters or merges in their backward lineage (Pignatelli et al., 2024).
-Finding #3: Base models like Llama-3.1-8B dominate the LLM supply chain, spawning thousands of derivatives, while task-specific models such as command-r-1-layer exhibit deep dependencies with other task-specific variants but avoid adapters or merges.
-4.4. RQ #4: Dataset Supply Chain Analysis
+Llama-3.1-8B 7,544 1,710 1,542 3,473 1,693 25
+Mistral-7B-v0.1 6,744 2,105 2,187 1,435 1,254 27
+Qwen2.5-7B 6,733 1,972 1,764 2,516 1,132 11
+Meta-Llama-3-8B 5,633 967 1,511 2,220 1,967 21
+Llama-3.1-70B 4,063 698 281 2,075 2,519 11
+Qwen2.5-32B 3,909 1,086 158 2,311 1,049 12
+Qwen2.5-1.5B 3,645 1,300 1,290 949 248 8
+Qwen2.5-0.5B 3,521 1,669 1,006 810 46 11
+Qwen2.5-14B 3,362 726 411 1,880 1,166 15
+Meta-Llama-3-8B-Instruct 3,118 640 405 1,394 1,305 34
 Table 4. Top-10 models sorted by backward subgraph size.
@@
 LeoPARD-0.8.1-4bit Quantization 16 1 14 15 DeepSeek-R1
-This research question aims to provide comprehensive insights into the different usages of the datasets within the LLM supply chain graph.
-There are two relationships between datasets.
-On one hand, one dataset can be created by combining a variety of other datasets from disparate sources.
-On the other hand, one dataset can serve as a building block for new datasets.
-In the LLM supply chain graph, 68,634 datasets act as training datasets that are used for model training or generating new datasets.
-Table 5 presents the top 10 datasets ranked by the highest number of included and derived datasets.
-(i) We observe that a training dataset can include many small datasets.
-Sitting at the top is macrocosm-os/images (Kido, 1961) from macrocosm.
-It is included in 550 datasets and is a general-purpose image modification corpus used in advanced multimodal AI research and large-scale vision-language model training.
-Moreover, bespokelabs/Bespoke (Elliot et al., 2024) from bespokelabs is part of 215 datasets, is a high-quality synthetic dataset designed to enhance multimodal AI research.
-Table 5. Top-10 datasets by # of included & derived datasets.
-Dataset # of included Dataset # of derived
-macrocosm-os/images 550 HuggingFaceH4 989
-bespokelabs/Bespoke 215 CodeFeedback 857
-databricks/databricks 139 MADLAD 850
-Open-Orca/OpenOrca 121 Capybara 823
-nguha/legalbench 117 Glot500 804
-OpenAssistant/oasst1 112 dolphin-coder 756
-LDJnr/Capybara 108 SlimOrca 726
-kvn420/Tenro_V4.1 105 orca_dpo_pairs 687
-teknium/OpenHermes 104 gutenberg 675
-Anthropic/hh-rlhf 98 samantha 658
-(ii) The other observation is that a single dataset can be derived to many datasets, showing the huge overlaps between the datasets and also the models trained with them.
-The right two columns of Table 5 show the top 10 datasets sorted by the maximum number of derived datasets they have been included in.
-The leading contributor, HuggingFaceH4 from HuggingFace has been derived to 989 datasets.
-It is primarily used for natural language understanding and generation tasks, including chatbots, text generation, code completion, reasoning, and multilingual processing, enabling the training and enhancement of models specialized in advanced conversational and instruction-following capabilities (Dietrich and Hollstein, 2025).
-Similarly, the CodeFeedback dataset from m-a-p has facilitated the creation of 857 datasets, which is a collection of high-quality code instruction queries designed to enhance the training of large language models (LLMs) for code generation, debugging, and explanation tasks, enabling improved performance in complex programming scenarios(Meyers et al., 2018).
-Finding #4: Datasets play critical roles in training with 68,634 datasets contributing to model development through inclusion and derivation, as seen in macrocosm-os/images and HuggingFaceH4.
-4.5. RQ #5: Supply Chain between Models and Datasets
-This research question aims to explore the interconnections between models and datasets within the supply chain graph.
-It provides valuable insights from dual perspectives, including one dataset versus multiple models and one model versus multiple datasets.
+Task-specific model analysis.
+Given a task-specific model, we want to understand how it evolves, e.g., what other models it relies on.
+To achieve that, for each model, we perform a backward analysis by running BFS following the incoming edges.
+To that end, the derived subgraph shows the models it relies on.
+Table 4 shows the top-10 task-specific models sorted by the backward subgraph size, which is the number of models they rely on.
+We make two interesting observations.
+(i) A fine-tuned model, command-r-1-layer, illustrates the depth and complexity of transformations in the LLM supply chain.
+This model operates in bfloat16 (BF16) precision for efficient text generation and natural language understanding (Pignatelli et al., 2024), originates from the base model c4ai, and has undergone extensive lineage evolution before reaching its final form.
+Specifically, it depends on 40 upstream artifacts, including 39 other fine-tuned models, and spans 39 transformation levels in its backward lineage chain, as detailed in Table 4.
+(ii) We observe that adapters are mainly used for lightweight fine-tuning and merges for model integration, but task-specific models like command-r-1-layer, as optimized standalone derivatives, do not evolve from adapters or merges in their backward lineage (Pignatelli et al., 2024).
+Finding #3: Base models like Llama-3.1-8B dominate the LLM supply chain, spawning thousands of derivatives, while task-specific models such as command-r-1-layer exhibit deep dependencies with other task-specific variants but avoid adapters or merges.
+4.4. RQ #4: Supply Chain of Models and Datasets
+This research question explores the interconnections between models and datasets from dual perspectives, including one dataset versus multiple models and one model versus multiple datasets.
 One dataset versus multiple models refers to the case when a single dataset is used to train multiple models.
-Table 6 shows the top 10 datasets based on the number of models trained on them.
-In particular, Mistral-7B-v0.1 takes the leading position and is a widely adopted open-source dataset known for its strong performance in general-purpose language understanding and generation tasks.
+Table 5 shows the top-10 datasets based on the number of models trained on them.
+In particular, (i) Mistral-v0.1 takes the leading position and is a widely adopted open-source dataset known for its strong performance in general-purpose language understanding and generation tasks.
 It has been used to train 1,093 models, including 300 fine-tuned variants, 300 adapters, 193 quantized models, and 300 merged models, highlighting its broad adoption across diverse model derivation strategies.
-In addition, the dataset TinyLlama-1.1B-v1.0, a compact and efficient model variant designed for low-resource deployment, is used to train 728 models, featuring 300 fine-tuned variants and 300 adapters.
-Similarly, open_llama_3b, an open-access dataset of LLaMA, supports 285 adapter-based models, indicating a preference for lightweight, modular adaptation.
-The dataset Yarn-Mistral-7b-128k also shows significant reuse, contributing to 279 adapters and 14 quantized models.
-Furthermore, MAmmoTH2-7B stands out with 210 merged models, showcasing its role in ensemble-style model fusion rather than traditional fine-tuning or adapter strategies.
-Lastly, the dataset Starling-LM-7B-alpha, known for alignment-focused training, contributes to a diverse range of downstream models, including 165 adapters and 18 quantized versions, illustrating its utility in fine-tuning and compression workflows.
-One model versus multiple datasets refers to the case when an LLM model is trained with multiple datasets.
-We observed that DeBERTa-ST-AllLayers-v3.1, a fine-tuned variant of the DeBERTa architecture, takes the top position, having been trained on 116 different datasets.
-Its adapter-based counterpart, DeBERTa-ST-AllLayers-v3.1bis, also leverages the same number of datasets via adapter-based training, emphasizing modular reuse across tasks.
 Table 5. Top-10 datasets sorted by # of models trained.
-Dataset Total Fine-tune Adapter Quantization Merges
-Mistral-7B-v0.1 1093 300 300 193 300
+Dataset Total Fine-tune Adapter Quantization Merge
+Mistral-v0.1 1,093 300 300 193 300
 TinyLlama-1.1B-v1.0 728 300 300 100 28
@@
 Starling-LM-7B-alpha 210 10 165 18 17
-In addition, models like static-similarity-mrl-mul-v1 and static-similarity-mrl-multilingual are both fine-tuned on 108 datasets, indicating their role in multilingual and multi-task similarity-based retrieval applications.
-The ModernBERT-base-embed and Llama-3.2-3B-Instruct families show strong dataset diversity as well, with 88 and 87 datasets, respectively, across fine-tuning and quantized variants (e.g., GGUF format).
-Interestingly, most of these models are fine-tuned, specifically, 8 out of the top 10, highlighting a trend where base models are transformed into task-specific variants through diverse training datasets.
-This pattern suggests that fine-tuning remains a dominant strategy for adapting base models to downstream tasks across heterogeneous data sources.
-Finding #5: Models and datasets exhibit strong bidirectional interdependence, with datasets like Mistral-7B-v0.1 spawning hundreds of models, while models such as DeBERTa-ST-AllLayers-v3.1 leverage diverse datasets to enhance adaptability, highlighting the critical role of dataset-model interactions in advancing AI.
-4.6. RQ #6: Dynamic Update Evaluation
-Figure 5. The number of changed models and datasets on Hugging Face from June 25 to July 15, 2025.
-This research question aims to understand the dynamic update of the LLM supply chain.
-Following the method discussed in Section 3.4, we perform a daily-based data collection by capturing how many nodes and edges are added and deleted each day.
-Figure 5 illustrates the sum of daily added and deleted of six key node categories (base, fine-tuned, adapters, quantized, merge variants, and datasets) from June 25 to July 15, 2025.
+(ii) The dataset TinyLlama-1.1B-v1.0 has been used to train 728 models, featuring 300 fine-tuned variants and 300 adapters.
+Similarly, open_llama_3b, an open-access dataset of Llama, supports 285 adapter-based models, indicating a preference for lightweight, modular adaptation.
+(iii) Further, MAmmoTH2-7B stands out with 210 merged models, showcasing its role in ensemble-style model fusion rather than traditional fine-tuning or adapter strategies.
+One model versus multiple datasets refers to the case when an LLM model is trained with multiple datasets.
+Table LABEL:tab:dataset_model_2 shows the top-10 models ranked by the number of datasets used for training.
+We observe that DeBERTa-ST-AllLayers-v3.1, a fine-tuned variant of the DeBERTa model, takes the top position, having been trained on 116 different datasets.
+Its adapter-based counterpart, DeBERTa-ST-AllLayers-v3.1bis, also leverages the same number of datasets via adapter-based training, emphasizing modular reuse across tasks.
+Table 6. Top-10 models sorted by # of training datasets.
+Model
+Model Type
+# of training dataset
+DeBERTa-ST-AllLayers-v3.1 Fine tune 116
+DeBERTa-ST-AllLayers-v3.1bis Adapters 116
+static-similarity-mrl-mul-v1 Fine tune 108
+static-similarity-mrl-multilingual Fine tune 108
+ModernBERT-base-embed Fine tune 88
+Llama-3.2-3B-Instruct Fine tune 87
+Llama-3.2-3B-Instruct-GGUF Quantization 87
+DavidLanz-3.2-3B-Instruct Fine tune 87
+static-retrieval-mrl-en-v1 Fine tune 79
+XLMRoBERTaM3-CustomPoolin Fine tune 72
+Figure 5. The number of changed datasets and models on Hugging Face from June 25 to July 15, 2025.
+In addition, models like static-similarity-mrl-mul-v1 and static-similarity-mrl-multilingual are both fine-tuned on 108 datasets, indicating their roles in multilingual and multi-task similarity-based retrieval applications.
+Interestingly, most of these models are fine-tuned, specifically, 8 out of the top-10, suggesting that fine-tuning remains a dominant strategy for adapting base models to downstream tasks across heterogeneous data sources.
+Finding #4: Models and datasets exhibit strong bidirectional interdependence, with datasets like Mistral-v0.1 spawning hundreds of models, while models such as DeBERTa-ST-AllLayers-v3.1 leverage different datasets to enhance adaptability, highlighting the critical roles of dataset-model interactions in advancing AI.
+4.5. RQ #5: Dynamic Update Evaluation
+This research question aims to understand the dynamic update of the LLM supply chain. We perform a daily-based data collection by capturing the addition and deletion of nodes and edges each day.
+Figure 5 denotes the number of changed (additions and deletions) datasets and models from June 25 to July 15, 2025.
 We make three interesting observations.
@@
 Noticeable spikes, such as on July 7 and July 9, align with major events like the Mistral-Fusion-v3 fine-tuning wave and dataset updates such as HFTime2025-News.
-(iii) Furthermore, adapter uploads peaked at 1,249 on June 28, while quantized variants reached 381 on July 9, driven by releases such as QWin-GGUF-7B.
+(iii) Furthermore, adapter uploads peaked at 1,249 on June 28, while quantized variants reached 381 on July 9, driven by releases such as Qwen-GGUF-7B.
 These patterns demonstrate HuggingGraph’s ability to capture evolving supply chain dynamics at a fine-grained level.
-Finding #6: The LLM supply chain exhibits continuous and high-volume daily changes, driven by frequent additions and deletions of models and datasets.
+Finding #5: The LLM supply chain exhibits continuous and high-volume daily changes, driven by frequent additions and deletions of models and datasets.
 This reflects a rapidly evolving and highly dynamic ecosystem shaped by active contributor behavior.
-4.7. RQ #7: Generability to Other AI Platforms
+4.6. RQ #6: Generalization to Other AI Platforms
 To validate HuggingGraph’s generalizability beyond Hugging Face, we applied our pipeline to another AI platform, Kaggle (Bojer and Meldgaard, 2021).
-As of July 25, 2025, Kaggle hosts 470 base and task-specific models with 3,146 variants and approximately 502K datasets.
+As of July 25, 2025, Kaggle hosts 470 base models, 3,146 task-specific models, and $\sim$502K datasets.
 Using Kaggle’s kernel and dataset APIs, we collected 2,640 models and 105,867 datasets.
-This significant gap is due to lack of the models’ and datasets’ metadata.
+This significant gap is due to a lack of models’ and datasets’ metadata.
 Of the datasets retrieved ($\sim$105K), only 137 datasets were included in our graph, as most lacked standardized documentation or traceable links to models.
 Many are standalone, poorly described, or lack contextual information, a challenge also observed on Hugging Face.
-We follow the same way to construct a heterogeneous graph consisting of 2,777 nodes, which include 2,640 model nodes, comprising 59 base models, 2,410 fine-tuned, 171 quantization models, and 137 datasets.
-The graph contains a total of 3,990 edges, on which we observed seven types of edges, (i) base model $\rightarrow$ fine-tuned model (467 edges), (i) base model $\rightarrow$ quantization model (62 edges), (iii) fine-tuned model $\rightarrow$ fine-tuned model (1,696 edges), (iv) fine-tuned model $\rightarrow$ quantized model (107 edges), (v) quantized model $\rightarrow$ quantized model (1 edges), (vi) dataset $\rightarrow$ fine-tuned model (1,614 edges), and (vii) dataset $\rightarrow$ quantization model (43 edges).
+We follow the same way to construct a heterogeneous graph consisting of 2,777 nodes, which include 2,640 model nodes, comprising 59 base models, 2,410 fine-tuned, 171 quantization models, and 137 dataset nodes.
+The graph contains a total of 3,990 edges, on which we observed seven types of edges, (i) base model $\rightarrow$ fine-tuned model (467 edges), (ii) base model $\rightarrow$ quantization model (62 edges), (iii) fine-tuned model $\rightarrow$ fine-tuned model (1,696 edges), (iv) fine-tuned model $\rightarrow$ quantized model (107 edges), (v) quantized model $\rightarrow$ quantized model (1 edge), (vi) dataset $\rightarrow$ fine-tuned model (1,614 edges), and (vii) dataset $\rightarrow$ quantization model (43 edges).
 We observed that the average degree is 1.44, indicating that the graph is sparse.
 We made two interesting observations.
-(i) The degree distribution is heavy-tailed and skewed: out of 2,777 total nodes, 2,305 nodes ($\approx$83%) have a total degree of 1.
+(i) The degree distribution is heavy-tailed and skewed: out of 2,777 total nodes, 2,305 nodes (83%) have a total degree of 1.
 Most nodes have low degrees, while a few highly connected hubs dominate the graph.
-For example, in-degrees range from 0 to 4 (with NaN peaking at 4), and tensorflow/mobilenet-v1 has the highest out-degree of 64, followed by google/nnlm with 56.
-(ii) Furthermore, the graph contains 448 weakly connected components (WCCs), reflecting high fragmentation.
+For example, in-degrees range from 0 to 4, and tensorflow/mobilenet-v1 has the highest out-degree of 64, followed by google/nnlm with 56.
+(ii) Furthermore, the graph contains 448 WCCs, reflecting high fragmentation.
 However, the largest WCC includes 65 nodes, suggesting the presence of a moderately sized core subgraph.
-Finding #7: The resulting graph exhibits structural properties consistent with our Hugging Face analysis, including a heavy-tailed degree distribution, sparse connectivity, and strong modular fragmentation, demonstrating the robustness and generalizability of our pipeline across platforms despite metadata limitations.
+Finding #6: The resulting graph exhibits structural properties consistent with our Hugging Face analysis, including a heavy-tailed degree distribution, sparse connectivity, and strong modular fragmentation, demonstrating the robustness and generalizability of our pipeline across platforms despite metadata limitations.
 5. Use Case
-HuggingGraph presents a technique to analyze the supply chain of the LLM ecosystem.
-The proposed graph can be used for various applications, e.g., auditing provenance, identifying biases, and revealing trends like quantized model scarcity, aiding researchers and platform maintainers.
-We discuss the following two use cases.
-Use case #1: Tracing lineage and dependencies in the LLM supply chain.
-In the LLM ecosystem, models are frequently built upon others through fine-tuning, adapter training, or quantization, forming complex chains of dependencies.
+HuggingGraph presents a technique to analyze the supply chain of the LLM ecosystem, which can be used for various applications, e.g., auditing provenance, identifying biases, and revealing trends like quantized model scarcity.
+We discuss the two use cases.
+Use case #1: Tracing lineage and dependency in LLM supply chain.
+Models are frequently built upon others through fine-tuning, adapter training, or quantization, forming complex chains of dependencies.
 However, when these relationships are not explicitly visible, it becomes difficult to verify where a model comes from, whether it inherits bias from upstream datasets, or if it complies with licensing constraints.
@@
 Use case #2: Identifying critical nodes and structural vulnerabilities.
-In the LLM ecosystem, certain models (e.g., gemma-2b) and datasets (e.g., The Pile) are reused so frequently that they become critical structural hubs, where failure or removal of them could disrupt numerous downstream dependencies.
+In the LLM ecosystem, certain models (e.g., gemma-2b) and datasets (e.g., The Pile) are reused so frequently that they become critical structural hubs, where their failure or removal could disrupt numerous downstream dependencies.
 These hidden single points of failure are difficult to detect without a comprehensive view of resource interconnections.
@@
 This visibility enables maintainers to safeguard vital assets and helps developers mitigate the risk of overreliance on fragile or under-maintained components.
-6. Related Work
-This section aims to provide in-depth coverage of the current research landscape concerning the opportunities within the LLM supply chain, mainly emphasizing the different dimensions of the relationship between the models and datasets, focusing on the model and dataset hosting platforms (Sindhu et al., 2024; Hadi et al., 2024).
+6. Related Work and Discussion
 LLM supply chain perspectives in AI.
-LLMs mark the revolutionary era in AI, with the edge passing through an exponential growth phase.
-A recent paper presents LLMs key components, including model infrastructure, lifecycle, and downstream applications (Wang et al., 2024).
-Another work showed that reusing these models has become widespread, which encourages the sharing and adaptation of base models on a larger scale (Jiang et al., 2023).
-An open-source AI ecosystem, such as Hugging Face, hosts a broad range of LLMs and datasets and, therefore, plays an important role in democratizing AI technologies (Riva et al., 2024).
-The base models are key ingredients in such an ecosystem (Xu et al., 2024b).
-They capture vast amounts of information from various datasets, allowing for effective task-specific variants like fine-tuning (Han et al., 2021).
-This trend underscores the significant impact of technological advancements on AI democratization and innovation (Cupać et al., 2024).
+LLMs are advancing across model infrastructure, lifecycle, and applications (Wang et al., 2024).
+Model reuse is widespread, promoting large-scale sharing and adaptation of base models (Jiang et al., 2023).
+Open-source ecosystems such as Hugging Face host diverse LLMs and datasets, democratizing AI (Riva et al., 2024).
+Base models (Xu et al., 2024b), trained on broad datasets, enable task-specific variants via fine-tuning (Han et al., 2021), reinforcing democratization and innovation (Cupać et al., 2024).
 Relationship analysis between LLM models and datasets.
-A recent study investigates the practical adaptation of base models to specific areas and tasks.
+A recent study investigates the practical adaptation of base models to specific tasks.
 Multitask fine-tuning has demonstrated the potential to enhance performance on target tasks with scarce labels (Xu et al., 2024a).
@@
 The Quadapter technique for language models tackles quantization difficulties by incorporating learnable parameters that scale activations channel-wise, mitigating overfitting during quantization-aware training (Park et al., 2022).
-These findings underscore the efficacy of various adaptation approaches in improving base model performance across areas.
-From the above, we can see that models heavily depend on each other for fine-tuning, adaptation, quantization or merger.
-Building on these insights, two other works have indicated that serious security vulnerabilities may encumber the LLM supply chain to a few developers (Kathikar et al., 2023; Osborne et al., 2024).
-These factors may affect the diversity of innovation.
-The authors relate several challenges arising during the software engineering, security, and privacy of different relations and components involved in model creation and deployment (Huang et al., 2023; Wang et al., 2024).
-Moreover, the vulnerabilities can be transferable from one model to another model during the fine-tuning process (Hajipour et al., 2024).
+In future, we would like to explore more through the following perspectives, (i) collecting more LLM supply chain information accurately and scalably; (ii) understanding the LLM supply chain better via both fundamental graph analytics (Feng et al., 2023; Hu et al., 2025; Li et al., 2025) and graph AI techniques (Kipf, 2016; Fu et al., 2022; Wang et al., 2025; Ji, 2025); and (iii) exploring the security and privacy threats on the LLM supply chain (Ji et al., 2021a; He et al., 2024a; Ji et al., 2021b; Cui et al., 2023; Brucker-Hahn et al., 2024).
 7. Conclusion
 This project studies the relationships between models and datasets in the LLM ecosystem, which are the central parts of the LLM supply chain.
-First, we design a methodology to systematically collect the supply chain information of LLMs.
-With that, we construct a large directed heterogeneous graph, having 402,654 nodes and 453,469 edges.
-Then, we perform different types of analysis and make multiple interesting findings.
+First, we systematically collect the supply chain information of LLMs.
+With that, we construct a directed heterogeneous graph, having 402,654 nodes and 462,524 edges.
+Lastly, we perform different types of analysis and make multiple interesting findings.
 Acknowledgment
@@
 Government.
+Disclaimer on Generative AI Usage
+This work does not use generative AI tools (e.g., ChatGPT, Copilot, Bard, Claude, etc.) for ideation, analysis, code development, or writing.
+All content, experiments, and results presented in this paper were created and validated solely by the authors.
+Mentions of generative AI are included only for scholarly context or comparison.